### v2

# Less is More: Improving LLM Alignment via Preference Data Selection

Abstract
Direct Preference Optimization (DPO) has emerged as a promising approach for aligning large language models with human preferences. While prior work mainly extends DPO from the aspect of the objective function, we instead improve DPO from the largely overlooked but critical aspect of data selection. Specifically, we address the issue of parameter shrinkage caused by noisy data by proposing a novel margin-maximization principle for dataset curation in DPO training. To accurately estimate margins for data selection, we propose a dual-margin guided approach that considers both external reward margins and implicit DPO reward margins. Extensive experiments demonstrate that our method reduces computational cost dramatically while improving performance. Remarkably, by using just 10% of the Ultrafeedback dataset, our approach achieves 3% to 8% improvements across various Llama and Mistral series models on the AlpacaEval 2.0 benchmark. Furthermore, our approach seamlessly extends to iterative DPO, yielding a roughly 3% improvement with 25% online data, while further reducing training time. These results highlight the potential of data selection strategies for advancing preference optimization.
Machine Learning, ICML
1 Introduction
Reinforcement Learning from Human Feedback (RLHF; Christiano et al., 2017; Ziegler et al., 2019) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values. Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization. However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
As an alternative, Rafailov et al. (2024) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data. DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF. Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024) and iterative learning schemes (Xiong et al., 2024). While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
This work thoroughly studies the impact of preference data quality on DPO training, which is crucial for developing more efficient training strategies. In particular, we achieve both improved performance and reduced computational costs through strategic data selection. Our research makes three primary contributions:
(1) We prove in theory the necessity of data selection in the presence of exogenous noise. Specifically, the inherent noise in the reward model may flip the preference between response pairs, leading to the emergence of the parameter shrinkage issue. Furthermore, we demonstrate that margin-based selection criteria can effectively address this issue by inducing parameter inflation.
(2) Building on this insight, we propose a margin-maximization principle for dataset curation that incorporates signals from the ensemble of external rewards and DPO implicit rewards. Through extensive experiments across diverse datasets and base models, we show that this selection strategy shows two consistent advantages: it substantially reduces computational overhead via efficient data selection and improves model performance compared to training on the full dataset. In particular, on the Ultrafeedback dataset and its variants, our method identifies a $10\%$ data subset for DPO training on LLama and Mistral series models, consistently achieving 3% to 8% point improvements on the AlpacaEval 2.0 benchmark relative to training on the complete dataset.
(3) Finally, we extend our data selection framework to iterative DPO settings, showing that selectively sampling online data can simultaneously lower computational costs and improve performance. In particular, we achieve 48.49% win rate and 54.99% length-control win rate on the AlpacaEval 2.0 benchmark using only $25\%$ of the online data for training.
Our findings provide both theoretical insights into the dynamics of preference learning and practical guidelines for more efficient DPO implementations. This work bridges an important gap between algorithmic innovation and data quality considerations in the context of LLM alignment.
2 Background
Reinforcement Learning from Human Feedback (RLHF) has emerged as a key method for aligning LLMs with human preferences. It leverages training data of the form $\mathcal{D}=\{x,y_{w},y_{l}\}$, where $x$ represents the input prompt, and $y_{w}$ and $y_{l}$ denote the preferred and dispreferred responses, respectively. The RLHF pipeline typically involves two stages: reward learning and policy optimization.
Reward Learning. In the reward learning stage, a reward model is trained to approximate human preferences based on preference data. By adopting the Bradley-Terry model (Bradley & Terry, 1952) to capture human preference, reward training involves minimizing the loss:
$\mathcal{L}_{\mathrm{RM}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\big{[}\log\sigma\big{(}r(x,y_{w})-r(x,y_{l})\big{)}\big{]},$
where $\sigma(\cdot)$ is the sigmoid function.
Policy Optimization with Reinforcement Learning. Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model. This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; Schulman et al., 2017) to optimize the policy by maximizing the expected reward.
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[r(x,y)-\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\right],$
where $\beta>0$ is the regularization parameter. However, this RL approach can be computationally expensive, sensitive to reward misspecification and require careful hyperparameter tuning.
Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; Rafailov et al., 2024) has been proposed. DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training. Instead of defining and optimizing a reward function explicitly, DPO minimizes
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right).$
By bypassing the intermediate step of reinforcement learning, DPO offers a more stable and computationally efficient alternative to standard RLHF, while still aligning models effectively with human feedback.
3 Methodology
We follow the model from Zhu et al. (2023) to illustrate why data selection can improve model performance. We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$. For reward learning, our reward model can be an explicit $r(x,y)$ (Ouyang et al., 2022), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (Rafailov et al., 2024). Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model. However, both humans and other large models may use inaccurate reward functions to generate labels, where the labels represent the ranking of two responses. We say the preference between $y_{w}$ and $y_{l}$ is generated by $r(x,y_{w})-r(x,y_{l})+\zeta$ where $\zeta$ is an exogenous error. We use $\Delta\phi(x)$ to denote $\phi(x,y_{w})-\phi(x,y_{l})$ for simplicity.
Parameter Shrinkage. Here, we hope to find $\omega$ to minimize
$\mathcal{L}_{\mathrm{RM}}(\omega)$ $=-\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\log(\frac{1}{1+e^{-\langle\Delta\phi(x),\omega\rangle}})$
$+\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\log(\frac{1}{1+e^{\langle\Delta\phi(x),\omega\rangle}})\Big{]}.$ (1)
It holds that the first-order condition is
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
$=$ $\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$ (2)
Since we know that $\langle\Delta\phi(x),\omega^{*}\rangle$ is positive, when $\zeta$ is small comparing to the margin, it holds that $\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}$ is convex with respect to $\zeta$. Due to Jensen’s inequality, it holds that
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
$\geq$ $\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
Similarly, we have
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
$\leq$ $\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
Since the optimal $\omega$ is $\omega^{*}$ without $\zeta$, plugging $\omega^{*}$ in Equation 2 will cause the left-hand side to be greater than the right-hand side. Therefore, the optimal $\omega$ with the existence of $\zeta$ intends to shrink to the original point compared to $\omega^{*}$ so that the first-order condition is still satisfied.
We provide the underlying intuition with an extreme example. If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Section 3. In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero. Thus, data selection is essential for acquiring policies with good performance. Finally, we remark that $\zeta$ can come from multiple resources, including human classification errors, different embeddings or reward models from other LLMs and so on.
Parameter Inflation. We next explain why selecting data points based on the margin can lead to parameter inflation, thereby offsetting the parameter shrinkage caused by errors.
First, when the margin is large, namely, $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$ is large, from the S-shaped graph of $\sigma(\cdot)$, we know that the slope is very small in this area. As a result, the probability of preference reversal caused by $\zeta$ is low, which means the likelihood of incorrect samples is also low.
Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side. Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Section 3.
Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Section 3 continuously. Hence, some techniques like regularization or early stopping when training are indispensable.
In summary, inaccuracies in the reward model can cause the parameters of LLMs to shrink toward zero. By selecting data with larger margins, we can compensate for the performance degradation caused by this shrinkage. The balance between parameter shrinkage and inflation offers the potential to enhance the performance of LLMs.
3.1 Dual-Margin guided Data Selection for Preference Learning
In this section, we present the Dual-Margin (DM) guided efficient preference data selection that enhances the performance of direct preference optimization while significantly reducing the training cost. Driven by the theoretical result, our main idea is to let the model reach an overall high margin during preference learning. We realize this by providing a robust estimation of reward margins for the entire dataset ahead of full-set training, which then allows efficient high-margin data filtering. DM operates in two stages: (1) computing external and implicit rewards, and (2) fusing these rewards to suppress noise and achieve more reliable margin estimates.
Initialization.
We aim select a subset $\mathcal{D}_{\mathrm{train}}$ from $\mathcal{D}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$, such that the model $\pi_{\theta}$ trained on the $\mathcal{D}_{\mathrm{train}}$ achieves performance comparable or better than that trained on $\mathcal{D}$. To guide this selection, we leverage an external reward model $r_{\mathrm{ex}}(x,y)$ and a small seed dataset $\mathcal{D}_{0}$ to learn an implicit reward model $r_{\mathrm{im}}(x,y)$.
Step 1: Reward calculation.
First, we calculate $r_{\mathrm{ex}}(x^{i},y_{w}^{i})$ and $r_{\mathrm{ex}}(x^{i},y_{l}^{i})$ for each datum in $\mathcal{D}$. Rafailov et al. (2024) proves that we can calculate implicit reward as
$r_{\mathrm{im}}(x,y)=\log\frac{\pi_{\theta}(y|x)}{{\pi_{\mathrm{ref}}}(y|x)}=\log\frac{\prod_{j}\pi_{\theta}(y_{j}|x,y_{1,\cdot,j-1})}{\prod_{j}\pi_{\mathrm{ref}}(y_{j}|x,y_{1,\cdot,j-1})}.$
As this reward margin between chosen and rejected samples calculated by different models shows a strong correlation (refer to Figure LABEL:fig:visual), we propose disentangling its calculation from the target model $\pi_{\theta}$. In particular, we introduce a small model $\pi_{s}$, and fine-tune it on $\mathcal{D}_{0}$ to get a weakly aligned model $\pi_{s}^{1}$. These two models are then utilized for low-cost implicit rewards calculation on $\mathcal{D}$.
Step 2: Margin fusion.
After calculating the margins $m_{\mathrm{ex}}=r_{\mathrm{ex}}(x,y_{w})-r_{\mathrm{ex}}(x,y_{l})$ and $m_{\mathrm{im}}=r_{\mathrm{im}}(x,y_{w})-r_{\mathrm{im}}(x,y_{l})$, we consider appropriate fusion strategies for combining these two signal sources to reduce noise in the reward modeling process. The simplest approach, which we call DM-ADD, directly sums the signals. This strategy reflects a lenient selection criterion where samples are considered valuable if they demonstrate a high margin from either reward source.
A strict fusion approach considers that samples should receive low priority if they show a low margin in either signal. We first transform margin values to margin-guided probability through a simple linear transformation
$\mathbb{P}(m)=\frac{\mathrm{clip}(m,M_{1},M_{2})-M_{1}}{M_{2}-M_{1}},\quad\text{ for }m\in\{m_{\mathrm{ex}},m_{\mathrm{im}}\},$
where $\mathrm{clip}(m)=\min(\max(m,M_{1}),M_{2}))$ and $(M_{1},M_{2})$ are tuning parameters. Following the fusion principle in multi-view probabilistic clustering (Deng et al., 2024), we further obtain
$\mathbb{P}(y_{w}\geq y_{l}|m_{\mathrm{ex}},m_{\mathrm{im}})$
$=$ $\frac{\mathbb{P}(m_{\mathrm{ex}})\mathbb{P}(m_{\mathrm{im}})}{\mathbb{P}(m_{\mathrm{ex}})\mathbb{P}(m_{\mathrm{im}})+(1-\mathbb{P}(m_{\mathrm{ex}}))\cdot(1-\mathbb{P}(m_{\mathrm{im}}))}.$ (3)
This adaptive approach mitigates the adverse effects of outlier samples with unusually high margin values. See more implementation details in Appendix B. Overall, this fusion approach is referred to as DM-MUL.
Sample Selection.
We select the samples with the highest fused margins to construct $\mathcal{D}_{\mathrm{train}}$. After selection, the subset can be used for DPO training on any target model.
3.2 Data Efficient Online Iterative DPO
Online RLHF (Xiong et al., 2024) has gained significant attention for its effectiveness in aligning models with human preferences. The method iteratively generates multiple responses for given prompts and employs an external reward model to identify Best-of-N (BoN) and Worst-of-N (WoN) responses, forming preference pairs. This ‘self-reward maximization’ approach effectively sharpens the model’s response space by reducing the probability of generating low-quality responses (Huang et al., 2024). However, the online generation process still produces ambiguous samples, particularly when prompts either have highly deterministic answer patterns or are exceptionally challenging for models to address. Such on-policy data pairs may not effectively guide model improvement. To address this problem, we propose a straightforward solution called DPO-DM: incorporating DM into each round of the on-policy data collection process and conducting preference optimization exclusively on selected high-margin samples. This approach can substantially reduce training costs, particularly when dealing with large prompt spaces.
4 Experiments
We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare DM with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of DM in enhancing comprehensive dialogue ability in Section 4.3; next, we explore the online setting in Section 4.4. lastly, we perform an ablation study for the offline preference data selection in Section 4.5.
4.1 Experimental Setup
Table 1: GPT4 win rates vs. chosen answers on three tasks. DPO is implemented on subsets selected using different filtering strategies. The supervised fintuned model of Llama-3.2-3B-base is used for initialization. Each strategy selected 2,000 samples for DPO training across TL;DR, HH, and Ultrafeedback. Bold and underlined numbers indicate the best and runner-up performances, respectively.
Filter Init Rand External Margin Implicit Margin IFD Margin DM Full
Strategy TOP MID BOT TOP MID BOT TOP MID BOT ADD MUL
TL;DR 34.00 46.50 66.25 42.00 22.00 30.75 43.00 19.75 1.75 41.25 55.25 63.75 83.25 36.75
HH 70.25 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 85.25 87.25 92.00
Ultrafeedback 61.49 82.86 91.18 73.29 25.84 89.81 77.14 37.02 72.05 83.60 54.53 91.55 90.68 80.99
Preference Datasets.
We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (Bai et al., 2022), and (3) UltraFeedback (Cui et al., 2024), which comprises quality-scored model responses across diverse prompts from multiple sources. To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (Meng et al., 2024). In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model. The scores are given by the PairRM (Jiang et al., 2023b) reward model. Statistics about these datasets are in Appendix A.
Models.
We perform sample selection and DPO training experiments across three model architectures: Llama-3.2-3B (MetaAI, 2024), Llama-3-8B (Dubey et al., 2024), and Mistral-7B (Jiang et al., 2023a), under Base and Instruct setup. For the base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets. For the instruct setup, we directly use Llama-3-8B-Instruct or Mistral-7B-Instruct-v0.2 as the start of DPO training. Concerning external reward model for margin calculation, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) which is the best reward model at this scale according to the RewardBench leadboard.
Evaluation.
Following Rafailov et al. (2024), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately. For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (Li et al., 2023b) as our evaluation benchmark, which consists of 805 diverse questions.  [^1] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (Ye et al., 2023), whereas AlpacaEval 2.0 employs GPT4-1106-preview. As the ground truth oracle is unavailable, we assess model performance through win rates against reference responses (written by humans or LLMs like GPT4). Specifically, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue. We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments. More details are presented in Appendix A.
Baselines.
For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (Xia et al., 2024)), (2) Conditional Perplexity (IFD) (Li et al., 2023a), which utilizes token probability to measure informativeness. We use the difference in IFD scores between chosen and rejected responses for preference data selection. (3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards. For (2) and (3), We segment the data into TOP (highest positive difference), MID (near-zero difference), and BOT (lowest negative difference) subsets. Specifically, previous work (Wu et al., 2024) argues that low-gap preference pairs are more informative (i.e., chosen and rejected answers are similar), here we use the IFD-MID strategy to quantify this scheme and call it Low-Gap. For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by Xiong et al. (2024); Dong et al. (2024) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
Implementation Details.
The SFT training of the base model is carried out for two epochs with a learning rate of $2\times 10^{-5}$. Sample packing (Tunstall et al., 2024) is employed to accelerate the training, and we use a block size of 4096. As for DPO training, we follow Rafailov et al. (2024) and used a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$. The DPO training process runs for two epochs using the AdamW optimizer. The learning rate follows a cosine schedule with a 0.1 warm-up ratio, and the global batch size is set to 32. For implicit reward calculation, we employ a Llama-3.2-3B SFT model and its DPO-tuned model, which is fine-tuned on 2,000 randomly selected samples from the complete dataset. More implementation details are presented in Appendix B.
Table 2: Performance comparison on AlpacaEval 2.0 using DPO-trained models with different 6,000-sample subsets (10% of full set). Both SFT and Instruct variants of Llama-3-8B were evaluated. LC and WR denote length-controlled and raw win rate, respectively. Bold number denotes the best-performing selected subset.
Dataset UltraFeedback Llama-UltraFeedback
Model Llama-3-Base (8B) Llama-3-Instruct (8B) Llama-3-Base (8B) Llama-3-Instruct (8B)
Metric LC (%) WR (%) LC (%) WR (%) LC (%) WR (%) LC (%) WR (%)
Init 9.61 6.69 22.92 22.57 9.61 6.69 22.92 22.57
Full 17.32 15.30 28.64 26.54 19.92 16.45 32.31 32.44
Random 12.33 10.96 22.74 24.59 11.58 9.51 31.51 31.92
Low-Gap 13.93 11.40 28.19 27.95 11.12 7.87 34.95 34.25
M-Ex 16.61 14.81 26.28 25.24 21.11 18.63 35.10 34.80
M-Im 19.33 17.80 29.71 29.44 18.88 16.25 33.71 32.92
DM-ADD 18.64 18.46 29.75 29.62 20.98 18.29 35.97 35.79
DM-MUL 19.53 19.09 28.30 28.39 21.67 20.01 36.36 36.47
4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
Firstly, we compare DM and baseline strategies on three widely-used preference datasets: TL;DR, HH, and UltraFeedback. Using a Llama-3.2-3B model as our base architecture, we evaluate different selection methods, each sampling 2,000 training examples for DPO training. We use AlpacaEval as the test sets of UltraFeedback as it better reveals the degree of improvement. The results, measured by GPT4 win rates, are presented in Table 1. We summarize the findings below.
  • DM consistently achieves (near) optimal win rates across all three tasks, while all baselines show weak results on at least one task. This demonstrates that our method is more robust to useless or toxic samples in diverse domains. For example, DM-MUL achieves 17% higher win rates to second most effective baselines on TL;DR task, this is because the DM-MUL assigns low selection priority so long as a low margin is observed from one source of reward.
  • More data in DPO training does not always yield better results - using just 2-5% of carefully selected data can surpass the performance achieved with the full dataset. This finding suggests the presence of ineffective samples in these tasks, as evidenced by the performance patterns of X-MID and X-BOT strategies (where X represents implicit margin). The results in Table 1 reveal that external, implicit, and IFD margins sometimes can identify toxic samples, as models trained on such data perform significantly worse than the original SFT model. This phenomenon underscores the critical importance of data filtering in DPO training. highlighting that training data should undergo a thorough quality assessment before being used in the training process.
  • Among all methods, only implicit margin-BOT consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit rewards into the DM strategy design. Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences. As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
4.2.1 Analysis
We analyze DM from two perspectives: the joint distribution of different sources of reward margins and the training dynamics for subsets with different margin properties. We utilize the UltraFeedback datasets for this analysis.
The left and middle panels of Figure LABEL:fig:visual illustrate the sample distribution across external-implicit and implicit-implicit margin dimensions on UltraFeedback. Several key observations emerge: (1) The correlation between implicit and external reward margins is notably weak. In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa. This pattern is consistent across other datasets (see Appendix C.2). This divergence highlights the distinct preference patterns captured by these two reward types. Supporting this observation, Table 1 demonstrates that neither margin type consistently outperforms the other in subset selection: external margins prove more effective for TL;DR, implicit margins show superior performance on HH, while both perform comparably on UltraFeedback. These findings underscore the importance of combining both reward types in an ensemble approach. (2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B), and the robust correlation is observed on other datasets as well (see Appendix C.2). This validates our design of using a separate small model for implicit reward margin calculations.
The right panel of Figure LABEL:fig:visual shows DPO training loss curves across different subset selection strategies. TOP/BOT selections converge quickly to low loss values, while MID selections maintain steady, higher loss levels. Surprisingly, BOT-selected samples learn as quickly as TOP-selected ones, suggesting that ‘bad’ preferences are also easy to grasp for LLM. The pattern is consistent across different datasets and models (see Appendix D.1 for more results concerning loss and margin curves on other datasets). While this observation might explain proposals that use absolute margin values for selection (Muldrew et al., 2024), Table 1 reveals that TOP and BOT samples ultimately produce opposing effects despite similar training dynamics. However, we observe a difference between BOT area samples selected by external and implicit margins, where Im-BOT exhibits training dynamics that are more similar to Im-TOP area samples in terms of loss reduction speed.
4.3 AlpacaEval 2.0 Win Rate Comparison
In this section, we conduct a detailed examination of the chat task to understand how data filtering enhances both DPO training efficiency and models’ versatile conversational abilities compared to GPT-4-turbo, representing a key application area for preference learning. We use both Llama-3-8B (base) and (instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, a benchmark widely adopted in preference learning studies. The results of DM and various baselines are shown in Table 2.
Table 3: Iterative DPO: AlpacaEval 2.0 benchmark results across three iterations of DPO training with UltraFeedback prompts.
Method Llama-3-Base (8B) Llama-3-Instruct (8B)
LC (%) WR (%) Len LC (%) WR (%) Len
DPO (r1) 17.64 13.36 1496 40.51 43.90 2173
DPO (r2) 23.06 22.45 1897 42.51 49.23 2366
DPO (r3) 29.03 30.86 2736 44.51 53.12 2860
DPO-DM (r1) 16.35 13.09 1624 42.20 45.74 2091
DPO-DM (r2) 23.79 24.17 1901 46.40 50.60 2316
DPO-DM (r3) 32.31 33.91 2565 48.49 54.99 2774
DM-ADD and DM-MUL consistently outperform full dataset DPO training. Table 2 shows models trained on DM-selected subsets (using only 10% of data) achieve 2-4% higher raw and LC win rates compared to full dataset training across different models (base and instruct) and datasets (offline and on-policy). These consistent improvements demonstrate DM’s robustness and effectiveness, validating the value of data filtering for DPO training. In contrast, all baseline methods demonstrate inconsistent performance across different datasets and models. We attribute this instability to ambiguous samples whose margins vary significantly across different reward models, suggesting that proper fusion of reward signals is necessary to mitigate such noise. Appendix D.2 presents additional training loss and margin curves, illustrating distinct training dynamics across different subsets, and the trend is similar to that on the 3B model as in Appendix D.1.
On-policy data primarily benefits its corresponding on-policy model. Llama-3-8B-Instruct shows approximately 8% higher win rates when trained on the on-policy Llama-UltraFeedback data compared to the off-policy UltraFeedback data, while Llama-3-8B-base demonstrates notably smaller improvements between these datasets. This phenomenon persists in both full and subset DPO training.
4.4 Online Iterative DPO
We explore the benefits of data selection in iterative DPO training using prompts from UltraFeedback as in Xiong et al. (2024). Our methodology employs DM-MUL to filter 5,000 samples from a pool of 20,000 generated samples per iteration. We evaluate both the base and instruct versions of Llama-3-8B, conducting single-epoch training rounds with hyperparameters $\beta=0.1$ and learning rate $5\times 10^{-7}$. Table 3 presents our experimental findings, from which we draw two key observations:
While Best-of-N and Worst-of-N preference pairs from a strong reward model are informative, data filtering is still important to DPO training. DPO-DM performs better for both base and instruct settings using only 25% of the data. This can be attributed to the presence of numerous ambiguous, low-margin samples in the constructed online preference dataset. Furthermore, standard iterative DPO shows increasing output length across training rounds, but data filtering mitigates this issue. By round three, DPO-DM demonstrates both superior performance and better output length control.
Table 4: Ablation study on model choice (Mistral-7B-Instruct-v-0.2). Each strategy selects a 6,000-sample subset.
Method LC (%) WR (%) Length
Initial 17.11 14.72 1676
Full 18.00 18.77 4189
Random 19.42 16.64 1621
Low-Gap 13.31 14.35 2525
DM-ADD 23.67 18.21 1527
DM-MUL 26.04 20.53 1568
4.5 Ablation Study
A critical aspect of dataset filtering methods is their generalization capability - specifically, how well the method (selected subset) transfers to new models, similar optimization algorithms, etc. From this perspective, we examine DM’s robustness and effectiveness across several dimensions.
Data filtering remains effective for new architecture and datasets. While our initial experiments focused on the Llama architecture and tokenizers, we investigated DM’s effectiveness when applied to fundamentally different model architectures. We evaluate the Mistral-7B-Instruct model using its corresponding Mistral-UltraFeedback dataset, maintaining consistent experimental parameters with those described in Section 4.3. The AlpacaEval 2.0 evaluation results are presented in Table 4. Our experiments show that DM-MUL significantly outperforms baseline methods while maintaining output lengths (i.e. around 1600) comparable to the initial instruct model. We observe that models trained on the full dataset tend to generate notably longer responses (i.e., 4189), primarily due to their learned behavior of incorporating numerous hyperlinks (e.g., https://www.ibdb.com/) in an attempt to enhance response credibility. Similarly, models trained on the Low-Gap selected subset exhibit increased response lengths (i.e., 2525) and follow a comparable pattern of including hyperlinks. In contrast, models trained on DM-selected subsets do not display this behavior. We consider the inclusion of hyperlinks without proper RAG (Retrieval-Augmented Generation) (Lewis et al., 2020) or website verification mechanisms to be problematic, as such references can be unreliable.
DM selected subsets are effective across other preference learning algorithms. We examine the effectiveness of our selected subsets across different DPO variants and compare the raw/LC win rates of models trained on full or selected subsets by IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024) and SLiC (Zhao et al., 2023). The results are presented in Figure LABEL:fig:abl_po. Our results demonstrate that DM-M consistently outperforms full-set training across all three algorithms, achieving particularly notable improvements in raw/LC win rates - over 12% - for the IPO algorithm. This advantage is maintained even across these preference learning algorithms with varying data efficiency (as measured by the performance gap between randomly selected 6,000 samples and the full dataset). These findings highlight the significant value of sample filtering for other preference learning. Additional results related to instruct model training can be found in Appendix D.4.
[…]
5 Related Work
Preference Learning Algorithms.Reinforcement Learning from Human Feedback has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (Ouyang et al., 2022). While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed. These include but not limited other RL-based training algorithms (Li et al., 2023c; Zhong et al., 2024), rejection sampling (Dong et al., 2023; Gulcehre et al., 2023), conditional supervised fine-tuning (Lu et al., 2022; Yang et al., 2024; Zhang et al., 2024a), and Direct Preference Optimization (Rafailov et al., 2024). Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance. Following the introduction of DPO, numerous works (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024; Tang et al., 2024; Han et al., 2024; Xu et al., 2024; Hong et al., 2024; Park et al., 2024) have attempted to improve its performance by modifying the DPO objective.
Data Selection in LLM Fine-tuning. Data selection is crucial in LLM post-training (Wang et al., 2024), primarily due to two key observations: LLM training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (Swayamdipta et al., 2020; Deng et al., 2023). Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (Cao et al., 2024; Li et al., 2024; Xia et al., 2024), often adapting active learning query strategies (Ren et al., 2021) to assess sample uncertainty and diversity. However, data efficiency in preference learning remains relatively unexplored. While some studies studied annotation cost reduction in preference dataset creation (Muldrew et al., 2024; Yasunaga et al., 2024), our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance. Furthermore, our method extends to iterative DPO (Xiong et al., 2024) and its variants (Zhang et al., 2024b), where training data is generated online.
6 Conclusion
Our research bridges the gap between algorithmic advances and data-focused approaches in Large Language Model (LLM) alignment by systematically examining how preference data quality affects Direct Preference Optimization (DPO). We address the issue of parameter shrinkage caused by noisy data and introduce a dual-margin-based strategy for selecting high-quality training examples. This approach not only improves model performance but also significantly increases computational efficiency. Our extensive experiments show that the method maintains or enhances performance while using just around 10% of the original training data, demonstrated through consistent improvements on the AlpacaEval2 benchmark. Additionally, our framework successfully extends to iterative DPO applications. These results emphasize the importance of careful data curation in developing better alignment techniques and provide practical guidelines for future research and implementation.
Appendix A Datasets and Evaluation Details
Data information.
The detailed information about the datasets used in the experiments is shown in Table 5. The test sets of TL;DR and HH are sampled from their original large testing pool, and we utilize prompts in AlpacaEval as the test sets for all models trained on UltraFeedback and its variants. In particular, results in Table 1 utilize Text-Davinci-003 generated answers as reference response as it can better reflect how models’ ability varies with different training data, and results in other tables all utilize GPT4-1106-preview generated answer as reference response (i.e., AlpacaEval 2.0).
Table 5: Statistical information about the training and evaluation datasets used in the experiments.
            Dataset             $|\mathcal{D}_{train}|$             $|\mathcal{D}_{test}|$             Type
            TL;DR             92858             400             Summarization
            HH             86372             400             Helpful & Harmless
            UltraFeedback             61135             -             Hybrid
            Llama-UltraFeedback             59876             -             Hybrid
            Mistral-UltraFeedback             59720             -             Hybrid
            AlpacaEval             -             805             Chat
Evaluation details.
Win rate judgments serve as a key evaluation metric across our experiments. For the TL;DR and HH, we slightly modify the evaluation prompts suggested by Rafailov et al. (2024) for the win rates judge. We use the original prompt designed for AlpacaEval and AlpacaEval 2.0 benchmark in all testing.
{spverbatim}
**TL;DR win rate prompt**: Which of the following summaries does a better job of summarizing the most important points in the given forum post, without including unimportant or irrelevant details? A good summary is both precise and concise.
Post: Post
Summary A: Summary_A
Summary B: Summary_B
FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why. SECOND, on a new line, state only ”A” or ”B” to indicate your choice. Your response should use the format: Comparison: ¡one-sentence comparison and explanation¿ Preferred: ¡”A” or ”B”¿
{spverbatim}
**HH win rate prompt**: For the following query to a chatbot, which response is more helpful or harmless? A helpful response provides useful information, while a harmless response avoids causing/suggesting harm or discomfort.
Query: query
Response A: res_a
Response B: res_b
FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless. SECOND, on a new line, state only ”A” or ”B” to indicate which response is more helpful or harmless. Your response should use the format: Comparison: ¡one-sentence comparison and explanation¿ More helpful or harmless: ¡”A” or ”B”¿
Appendix B More Implementation Details
We present the implementation details of our baseline methods: TOP, MID, and BOT using margins calculated from three metrics, Conditional Perplexity (CPPL), External (Ex), and Implicit (Im) rewards. Subsequently, we describe the implementation of DM-MUL in our experimental setup.
The baseline strategies are implemented as follows: first, we eliminate outlier samples with extreme margin values (both positively high and negatively low) for CPPL, Ex, and Im metrics. For the TOP and BOT strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively. The MID strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
For DM-MUL, we set $M_{1}=-2$ as the lower bound for both external and implicit reward margins. The upper bound $M_{2}$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[M_{2},\text{max margin}]$ is less than 30, or (2) The number of samples in $[M_{2},\text{max margin}]$ is less than max $\text{max margin}-M_{2}$.
Appendix C Visualization of Margin Distributions
C.1 Singular Margin Distribution
The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures LABEL:fig:cppl-margin,LABEL:fig:ex-margin,LABEL:fig:im-margin, reveal a notable concentration of sample margins around zero. This clustering around the zero indicates ambiguous preference labels. It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 4.2.1.
C.2 Joint Margin Distribution
To complement the left and middle subfigures in Figure LABEL:fig:visual, we present additional results showing the joint margin distributions of samples on the other datasets in Figure LABEL:fig:visual-extra. Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation. These further enhance the rationality of our design detail of DM: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
Appendix D More Experimental Results
D.1 Train Loss and Margin Curves - 3B
To complement the right subfigure in Figure LABEL:fig:visual, we present additional results showing the progression of training loss and margins throughout the DPO training process. The results are shown in Figure LABEL:fig:loss-margin. All strategies demonstrated consistent patterns in loss reduction: both TOP margin-oriented and BOT strategies achieved rapid decreases in training loss, while the MID strategy exhibited slower convergence and remained at significantly higher final loss values. Regarding training margins, TOP strategies achieved higher levels compared to BOT and MID approaches. Notably, our proposed DM-ADD and DM-MUL strategies demonstrated even larger margins than the Implicit Margin-TOP strategy.
D.2 Train Loss and Margin Curves - 8B
Figure LABEL:fig:loss-margin-8b presents the progression of training loss and margins during DPO training of the 8B model. We show the fine-tuned version of the Llama-3-8B (base) model trained on UltraFeedback as a representative case since all models exhibited similar trends, as in Figure LABEL:fig:loss-margin.
D.3 Resources and computation cost
For all experiments, we utilized 8 A100 GPUs. We conduct SFT/DPO training with 4 A100 GPUs for all runs in our experiments. For both Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) training, we allocated 4 A100 GPUs per run. Training 8B parameter models on the UltraFeedback dataset for two epochs required approximately 9 hours of computation time. In each round of iterative DPO implementation, we performed generation and annotation processes on 4 A100 GPUs, with each GPU processing 5,000 prompts with 5 distinct generations per prompt. The overall generation that utilizes vLLM (Kwon et al., 2023) for acceleration takes about 1.5 hours, and the corresponding reward annotation takes about 2 hours.
D.4 More Results for Ablation Study on the DPO Variants
As a complementary study to the results shown in Figure LABEL:fig:abl_po, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters. The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by DM-MUL achieved significantly higher win rates across most evaluation scenarios.
D.5 Ablation Study on Hyper-parameters
We investigate the impact of hyperparameters $\beta$ and training epochs on DPO performance. Specifically, we train the Llama-UltraFeedback model using DPO with $\beta=0.01$ for both one and two epochs. The model’s performance is evaluated using AlpacaEval 2.0 win rates, with results shown in Table 6. Our experimental results demonstrate that DM-MUL selected subsets significantly enhance the performance of vanilla DPO on these different hyperparameter configurations. Notably, DPO-DM-MUL achieves superior results compared to state-of-the-art preference learning algorithms such as SimPO (Meng et al., 2024), despite the latter utilizing the complete dataset for DPO training.
Table 6: Ablation study on $\beta$ ($\beta=0.01$): performance comparison of methods trained on Llama-Ultrafeedback dataset, benchmarked against state-of-the-art algorithms trained on the complete dataset.
Method LC (%) WR (%) Epochs
DPO 38.65 39.57 2
DPO 40.30 37.90 1
DPO-DM-MUL 43.66 42.52 2
DPO-DM-MUL 45.17 41.86 1
KTO 33.10 31.80 1
IPO 35.60 35.60 1
SimPO 43.72 39.88 1
However, we observe that selecting a small $\beta$ value is not consistently a good choice for DPO training. Specifically, a small $\beta$ corresponds to a relaxed Kullback-Leibler (KL) divergence constraint in the policy optimization process. This relaxation can permit excessive deviation from the initial policy, potentially compromising the model’s learned behaviors and stability during training. For instance, when we conduct DPO training with Llama-3-8B-Instruct/Mistral-7B-Instruct-v0.2 on the Ex/Im-TOP selected 6,000 subsets from UltraFeedback, with $\beta=0.01$ and two epochs update, we find that although the model could respond normally to most questions, it sometimes outputs repeated or chaotic tokens, as shown in Figure 9. And their win rates on AlpacaEval 2.0 dropped by more than 10 points as a consequence.
Further analysis of the training details revealed a significant degradation in log probabilities for both chosen and rejected samples, coinciding with the model’s performance decline. Specifically, during the above-mentioned Mistral model DPO training, the log probability values for chosen samples decreased from -400 to -1400, while rejected samples showed a more dramatic reduction from -600 to -4600. Overall, Mistral suffers more from this log probability drop compared to Llama.
Such phenomenon can be avoided by using a smaller learning rate: from $5\times 10^{-7}$ to $3\times 10^{-7}$ or early stop at the end of epoch 1. These operations can lead to a relatively smaller drop in chosen/rejected log probability. However, we still do not recommend using small $\beta$, as more potential sacrifices in other tasks such as reasoning may also appear in this case.
Figure 9: The model breaking pattern when conducting DPO training with small $\beta$ ($\beta=0.01$) for Llama-3-8B-Instruct and Mistral-7B-Instruct-V0.2. We select two examples of abnormal responses given by each model.

### v3

@@
 Specifically, we address the issue of parameter shrinkage caused by noisy data by proposing a novel margin-maximization principle for dataset curation in DPO training.
-To accurately estimate margins for data selection, we propose a dual-margin guided approach that considers both external reward margins and implicit DPO reward margins.
-Extensive experiments demonstrate that our method reduces computational cost dramatically while improving performance.
-Remarkably, by using just 10% of the Ultrafeedback dataset, our approach achieves 3% to 8% improvements across various Llama and Mistral series models on the AlpacaEval 2.0 benchmark.
-Furthermore, our approach seamlessly extends to iterative DPO, yielding a roughly 3% improvement with 25% online data, while further reducing training time.
+To further mitigate the noise in different reward models, we propose a Bayesian Aggregation approach that unifies multiple margin sources (external and implicit) into a single preference probability.
+Extensive experiments in diverse settings demonstrate the consistently high data efficiency of our approach.
+Remarkably, by using just 10% of the Ultrafeedback dataset, our approach achieves 3% to 8% improvements across various Llama, Mistral, and Qwen models on the AlpacaEval2 benchmark.
+Furthermore, our approach seamlessly extends to iterative DPO, yielding a roughly 3% improvement with 25% online data, revealing the high redundancy in this presumed high-quality data construction manner.
 These results highlight the potential of data selection strategies for advancing preference optimization.
-Machine Learning, ICML
 1 Introduction
-Reinforcement Learning from Human Feedback (RLHF; Christiano et al., 2017; Ziegler et al., 2019) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values.
+Reinforcement Learning from Human Feedback (RLHF; christiano2017deep, ; ziegler2019fine, ) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values.
 Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization.
 However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
-As an alternative, Rafailov et al.
-(2024) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data.
+As an alternative, (rafailov2024direct, ) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data.
 DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF.
-Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024) and iterative learning schemes (Xiong et al., 2024).
+Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ) and iterative learning schemes (xiong2024iterative, ).
 While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
@@
 (1) We prove in theory the necessity of data selection in the presence of exogenous noise.
-Specifically, the inherent noise in the reward model may flip the preference between response pairs, leading to the emergence of the parameter shrinkage issue.
+Specifically, the noise in the reward model may flip the preference between response pairs, leading to the emergence of the parameter shrinkage issue.
 Furthermore, we demonstrate that margin-based selection criteria can effectively address this issue by inducing parameter inflation.
-(2) Building on this insight, we propose a margin-maximization principle for dataset curation that incorporates signals from the ensemble of external rewards and DPO implicit rewards.
-Through extensive experiments across diverse datasets and base models, we show that this selection strategy shows two consistent advantages: it substantially reduces computational overhead via efficient data selection and improves model performance compared to training on the full dataset.
-In particular, on the Ultrafeedback dataset and its variants, our method identifies a $10\%$ data subset for DPO training on LLama and Mistral series models, consistently achieving 3% to 8% point improvements on the AlpacaEval 2.0 benchmark relative to training on the complete dataset.
+(2) Driven by the theoretical results and the derived margin-maximization principle, we propose a Bayesian Aggregation for Preference data Selection (BeeS) strategy.
+BeeS incorporates signals from both external rewards and DPO implicit rewards, and deprioritizes a preference pair if it exhibits a low reward margin from any single reward source to mitigate potential noise.
+Through extensive experiments across diverse datasets and models, we show that this selection strategy shows two consistent advantages: it substantially reduces computational overhead via efficient data selection and improves model performance compared to training on the full dataset.
+In particular, on the UltraFeedback dataset and its variants, our method identifies a $10\%$ data subset for DPO training on LLama, Mistral, and Qwen series models, consistently achieving 3% to 8% point improvements on the AlpacaEval 2.0 benchmark relative to training on the complete dataset.
 (3) Finally, we extend our data selection framework to iterative DPO settings, showing that selectively sampling online data can simultaneously lower computational costs and improve performance.
@@
 This work bridges an important gap between algorithmic innovation and data quality considerations in the context of LLM alignment.
+1.1 Related Work
+Preference learning algorithms.
+Reinforcement Learning from Human Feedback also known as dueling RL (pacchiano2021dueling, ) or preference-based RL (chen2022human, ), has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (ouyang2022training, ).
+While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed.
+These include but not limited other RL-based training algorithms (li2023remax, ; zhong2024dpo, ), rejection sampling (dong2023raft, ; gulcehre2023reinforced, ), conditional supervised fine-tuning (lu2022quark, ; yang2024rewards, ; zhang2024reward, ), and Direct Preference Optimization (rafailov2024direct, ).
+Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance.
+Following the introduction of DPO, numerous works (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ; meng2024simpo, ; tang2024generalized, ; han2024f, ; xu2024contrastive, ; hong2024orpo, ; wu2024self, ) have attempted to improve its performance by modifying the DPO objective.
+Data selection in LLM Fine-tuning.
+Data selection is crucial in LLM post-training (wang2024survey, ) for two key observations: post-training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (swayamdipta2020dataset, ; deng2023counterfactual, ).
+Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (cao2023instruction, ; li2024superfiltering, ; xia2024less, ), often adapting active learning query strategies (ren2021survey, ) to assess sample uncertainty and diversity.
+However, data efficiency in preference learning remains relatively unexplored.
+Prior studies have studied reducing annotation costs in preference dataset creation (muldrew2024active, ; yasunaga2024alma, ) and on scenarios involving numerous ranking annotations (thekumparampil2024comparing, ; mukherjee2024optimal, ).
+Other research aims to improve a model’s ability to distinguish between two responses by adding a margin to the loss term (meng2024simpo, ; park2024disentangling, ; azar2024general, ).
+Additionally, concurrent work highlights the importance of margins for preference data filtering, though there is debate on whether hard samples help or hinder preference learning (wu2024beta, ; khaki2024rs, ; yu2025rip, ; gao2025principled, ).
+Our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance.
+Furthermore, our method extends to iterative DPO (xiong2024iterative, ) and its variants (zhang2024self, ), wherein training data is dynamically generated by the model during its iterative training process.
 2 Background
@@
 In the reward learning stage, a reward model is trained to approximate human preferences based on preference data.
-By adopting the Bradley-Terry model (Bradley & Terry, 1952) to capture human preference, reward training involves minimizing the loss:
+By adopting the Bradley-Terry model (bradley1952rank, ) to capture human preference, reward training involves minimizing the loss:
 $\mathcal{L}_{\mathrm{RM}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\big{[}\log\sigma\big{(}r(x,y_{w})-r(x,y_{l})\big{)}\big{]},$
@@
 Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model.
-This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; Schulman et al., 2017) to optimize the policy by maximizing the expected reward.
+This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; schulman2017proximal, ) to optimize the policy by maximizing the expected reward.
 $\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[r(x,y)-\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\right],$
 where $\beta>0$ is the regularization parameter. However, this RL approach can be computationally expensive, sensitive to reward misspecification and require careful hyperparameter tuning.
-Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; Rafailov et al., 2024) has been proposed.
+Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; rafailov2024direct, ) has been proposed.
 DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training.
 Instead of defining and optimizing a reward function explicitly, DPO minimizes
-$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right).$
+$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right).$
 By bypassing the intermediate step of reinforcement learning, DPO offers a more stable and computationally efficient alternative to standard RLHF, while still aligning models effectively with human feedback.
 3 Methodology
-We follow the model from Zhu et al.
-(2023) to illustrate why data selection can improve model performance.
+3.1 Parameter Shrinkage and Inflation Analysis
+We follow the model from (zhu2023principled, ) to illustrate why data selection can improve model performance.
 We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$.
-For reward learning, our reward model can be an explicit $r(x,y)$ (Ouyang et al., 2022), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (Rafailov et al., 2024).
+For reward learning, our reward model can be an explicit $r(x,y)$ (ouyang2022training, ), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (rafailov2024direct, ).
 Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model.
@@
 Parameter Shrinkage. Here, we hope to find $\omega$ to minimize
-$\mathcal{L}_{\mathrm{RM}}(\omega)$ $=-\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\log(\frac{1}{1+e^{-\langle\Delta\phi(x),\omega\rangle}})$
-$+\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\log(\frac{1}{1+e^{\langle\Delta\phi(x),\omega\rangle}})\Big{]}.$ (1)
+$\mathcal{L}_{\mathrm{RM}}(\omega)=-\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\log(\frac{1}{1+e^{-\langle\Delta\phi(x),\omega\rangle}})+\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\log(\frac{1}{1+e^{\langle\Delta\phi(x),\omega\rangle}})\Big{]}.$ (1)
 It holds that the first-order condition is
-$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
-$=$ $\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$ (2)
+$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}=\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$ (2)
 Since we know that $\langle\Delta\phi(x),\omega^{*}\rangle$ is positive, when $\zeta$ is small comparing to the margin, it holds that $\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}$ is convex with respect to $\zeta$.
 Due to Jensen’s inequality, it holds that
-$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
-$\geq$ $\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
+$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}\geq\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
 Similarly, we have
-$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
-$\leq$ $\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
+$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}\leq\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
 Since the optimal $\omega$ is $\omega^{*}$ without $\zeta$, plugging $\omega^{*}$ in Equation 2 will cause the left-hand side to be greater than the right-hand side.
@@
 We provide the underlying intuition with an extreme example.
-If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Section 3.
+If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Equation 1.
 In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero.
@@
 Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side.
-Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Section 3.
-Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Section 3 continuously.
+Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Equation 1.
+Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Equation 1 continuously.
 Hence, some techniques like regularization or early stopping when training are indispensable.
@@
 The balance between parameter shrinkage and inflation offers the potential to enhance the performance of LLMs.
-3.1 Dual-Margin guided Data Selection for Preference Learning
-In this section, we present the Dual-Margin (DM) guided efficient preference data selection that enhances the performance of direct preference optimization while significantly reducing the training cost.
-Driven by the theoretical result, our main idea is to let the model reach an overall high margin during preference learning.
+Driven by this theoretical result, our main idea is to let the model reach an overall high margin during preference learning.
 We realize this by providing a robust estimation of reward margins for the entire dataset ahead of full-set training, which then allows efficient high-margin data filtering.
-DM operates in two stages: (1) computing external and implicit rewards, and (2) fusing these rewards to suppress noise and achieve more reliable margin estimates.
-Initialization.
-We aim select a subset $\mathcal{D}_{\mathrm{train}}$ from $\mathcal{D}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$, such that the model $\pi_{\theta}$ trained on the $\mathcal{D}_{\mathrm{train}}$ achieves performance comparable or better than that trained on $\mathcal{D}$.
-To guide this selection, we leverage an external reward model $r_{\mathrm{ex}}(x,y)$ and a small seed dataset $\mathcal{D}_{0}$ to learn an implicit reward model $r_{\mathrm{im}}(x,y)$.
-Step 1: Reward calculation.
-First, we calculate $r_{\mathrm{ex}}(x^{i},y_{w}^{i})$ and $r_{\mathrm{ex}}(x^{i},y_{l}^{i})$ for each datum in $\mathcal{D}$. Rafailov et al. (2024) proves that we can calculate implicit reward as
-$r_{\mathrm{im}}(x,y)=\log\frac{\pi_{\theta}(y|x)}{{\pi_{\mathrm{ref}}}(y|x)}=\log\frac{\prod_{j}\pi_{\theta}(y_{j}|x,y_{1,\cdot,j-1})}{\prod_{j}\pi_{\mathrm{ref}}(y_{j}|x,y_{1,\cdot,j-1})}.$
-As this reward margin between chosen and rejected samples calculated by different models shows a strong correlation (refer to Figure LABEL:fig:visual), we propose disentangling its calculation from the target model $\pi_{\theta}$.
-In particular, we introduce a small model $\pi_{s}$, and fine-tune it on $\mathcal{D}_{0}$ to get a weakly aligned model $\pi_{s}^{1}$.
-These two models are then utilized for low-cost implicit rewards calculation on $\mathcal{D}$.
-Step 2: Margin fusion.
-After calculating the margins $m_{\mathrm{ex}}=r_{\mathrm{ex}}(x,y_{w})-r_{\mathrm{ex}}(x,y_{l})$ and $m_{\mathrm{im}}=r_{\mathrm{im}}(x,y_{w})-r_{\mathrm{im}}(x,y_{l})$, we consider appropriate fusion strategies for combining these two signal sources to reduce noise in the reward modeling process.
-The simplest approach, which we call DM-ADD, directly sums the signals.
-This strategy reflects a lenient selection criterion where samples are considered valuable if they demonstrate a high margin from either reward source.
-A strict fusion approach considers that samples should receive low priority if they show a low margin in either signal.
-We first transform margin values to margin-guided probability through a simple linear transformation
-$\mathbb{P}(m)=\frac{\mathrm{clip}(m,M_{1},M_{2})-M_{1}}{M_{2}-M_{1}},\quad\text{ for }m\in\{m_{\mathrm{ex}},m_{\mathrm{im}}\},$
-where $\mathrm{clip}(m)=\min(\max(m,M_{1}),M_{2}))$ and $(M_{1},M_{2})$ are tuning parameters.
-Following the fusion principle in multi-view probabilistic clustering (Deng et al., 2024), we further obtain
-$\mathbb{P}(y_{w}\geq y_{l}|m_{\mathrm{ex}},m_{\mathrm{im}})$
-$=$ $\frac{\mathbb{P}(m_{\mathrm{ex}})\mathbb{P}(m_{\mathrm{im}})}{\mathbb{P}(m_{\mathrm{ex}})\mathbb{P}(m_{\mathrm{im}})+(1-\mathbb{P}(m_{\mathrm{ex}}))\cdot(1-\mathbb{P}(m_{\mathrm{im}}))}.$ (3)
+3.2 Multi-source Margin Aggregation
+Building on our previous analysis, we aim to develop a data selection strategy based on the margin-maximization principle, with the calculation of reward margin being the critical component.
+We examine two distinct types of reward margin calculations: external reward margin and implicit reward margin.
+The external reward margin is provided by an external reward function, while the implicit reward margin is derived from the implicit reward $\log\frac{\pi_{\theta}}{\pi_{\mathrm{ref}}}$, where $\pi_{\theta}$ represents the policy trained by DPO (rafailov2024direct, ) (see Section 4.1 for details on margin calculation).
+We visualize the joint distribution of different reward margin sources using the UltraFeedback dataset (cui2023ultrafeedback, ) in Figure LABEL:fig:visual.
+  • The left and middle panels of Figure LABEL:fig:visual reveal several key phenomena: (1) The correlation between implicit and external reward margins is notably weak.
+In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa.
+This divergence highlights the distinct preference patterns captured by these two reward types.
+This underscores the need to combine both reward types for a reliable margin estimation.
+(2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B).
+Notably, these two patterns remain consistent across other datasets as well (see Appendix B.2).
+  • Online RLHF (xiong2024iterative, ; dong2025self, ) employs the target model to generate multiple responses for given prompts iteratively and uses an external reward model to identify the response pair with the largest margin for DPO training.
+The right panel of Figure LABEL:fig:visual illustrates that a max-margin pair construction method, even when derived from one strong reward model, can still yield ambiguous preferences when evaluated by another reward model (which shows similar performance on RewardBench Leadboard).
+This ambiguity in preference signal indicates that the online data generation process can still cause high redundancy, which may offer little to no benefit, or could even be detrimental, to online-DPO training.
+These observations highlight the need for multi-source margin aggregation to achieve a more robust margin estimation, thereby enhancing data selection and preference learning.
+To this end, we propose a strict aggregation strategy, Bayesian Aggregation for Preference data Selection (BeeS), that deprioritizes a preference pair if it exhibits a low reward margin from any single reward source.
+We implement this method through a general three-step procedure.
+Step 1: in-distribution pre-DPO training.
+Our objective is to obtain the in-distribution implicit reward model with low computational and GPU-memory cost.
+Given the strong correlation of that margin across different models (See Figure LABEL:fig:visual), this weak-to-strong guidance is feasible.
+To achieve this, we randomly select a small seed dataset $\mathcal{D}_{0}$ from $\mathcal{D}$ and employ a (or several) small model to perform preference learning on this seed set.
+The high sample and training efficiency of the DPO loss (kim2025spread, ) ensures the feasibility of this approach.
+Step 2: margin calculation.
+We calculate external and implicit reward margins as $m_{\mathrm{ex}}=r_{\mathrm{ex}}(x^{i},y_{w}^{i})-r_{\mathrm{ex}}(x^{i},y_{l}^{i})$ and $m_{\mathrm{im}}=r_{\mathrm{im}}(x^{i},y_{w}^{i})-r_{\mathrm{im}}(x^{i},y_{l}^{i})$ for each datum in $\mathcal{D}$, where we directly calculate $r_{\mathrm{im}}(x,y)$ as $\log\frac{\pi_{\theta}(y|x)}{{\pi_{\mathrm{ref}}}(y|x)}$.
+Here, $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$ denote the small model before and after preference learning.
+We assume that there are $K$ margins involved from these two typical sources.
+Step 3: Bayesian aggregation.
+To mitigate noise from individual reward sources, we utilize Bayesian probability theory for their robust aggregation.
+Given that reward margins from different sources often vary in their underlying distributions and value ranges, we propose projecting these diverse margins into a unified probability space.
+This projection serves to quantify the confidence that a specific preference direction $y_{w}>y_{l}$ is correct, formally expressed as $\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})$.
+Assuming that these sources are conditionally independent, then following the Bayesian formula transformation in previous work (liu2022mpc, ; deng2024a3s, ), the preference probability can be expressed as:
+$\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})=\frac{\prod_{i=1}^{K}\mathbb{P}(y_{w}>y_{l}|m^{i})}{\prod_{i=1}^{K}\mathbb{P}(y_{w}>y_{l}|m^{i})+\prod_{i=1}^{K}(1-\mathbb{P}(y_{w}>y_{l}|m^{i}))}.$ (3)
+The typical absence of well-defined preference labels (e.g., $y_{w}>y_{l}$ for clear preference, or $y_{w}=y_{l}$ for indifference) renders the rigorous estimation of single-source preference probabilities challenging.
+[^1] This estimation typically relies on methods such as isotonic regression or histogram analysis.
+Consequently, we approximate the probability using a linear projection: $\mathbb{P}(y_{w}>y_{l}|m^{i})=\frac{\mathrm{clip}(m^{i},L,U)-L}{U-L}$ , where $\mathrm{clip}(m,L,U)=\min(\max(m,L),U))$ and $L$, $U$ are tuning parameters.
 This adaptive approach mitigates the adverse effects of outlier samples with unusually high margin values.
-See more implementation details in Appendix B.
-Overall, this fusion approach is referred to as DM-MUL.
+See more implementation details and discussion about the derivation approximation of Eq.
+(3) in Appendix A.1.
 Sample selection.
-We select the samples with the highest fused margins to construct $\mathcal{D}_{\mathrm{train}}$. After selection, the subset can be used for DPO training on any target model.
-3.2 Data Efficient Online Iterative DPO
-Online RLHF (Xiong et al., 2024) has gained significant attention for its effectiveness in aligning models with human preferences.
-The method iteratively generates multiple responses for given prompts and employs an external reward model to identify Best-of-N (BoN) and Worst-of-N (WoN) responses, forming preference pairs.
-This ‘self-reward maximization’ approach effectively sharpens the model’s response space by reducing the probability of generating low-quality responses (Huang et al., 2024).
-However, the online generation process still produces ambiguous samples, particularly when prompts either have highly deterministic answer patterns or are exceptionally challenging for models to address.
-Such on-policy data pairs may not effectively guide model improvement.
-To address this problem, we propose a straightforward solution called DPO-DM: incorporating DM into each round of the on-policy data collection process and conducting preference optimization exclusively on selected high-margin samples.
-This approach can substantially reduce training costs, particularly when dealing with large prompt spaces.
+We consider the data selection for both the one-pass DPO training and the iterative DPO workflow.
+For the former, we directly select the samples with the highest aggregated preference probability to construct $\mathcal{D}_{\mathrm{train}}$.
+The threshold depends on how many preference samples we prefer to use (but should guarantee that samples with negative margins are excluded).
+For the latter, we only need to train implicit reward models in the first iteration, and BeeS three-step procedures are applied for each iteration.
 4 Experiments
-We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare DM with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of DM in enhancing comprehensive dialogue ability in Section 4.3; next, we explore the online setting in Section 4.4.
-lastly, we perform an ablation study for the offline preference data selection in Section 4.5.
+Table 1: GPT-4 judged win rates for Llama-3.2-3B models fine-tuned with DPO on subsets selected by various data selection strategies.
+For every strategy and benchmark (TL;DR, HH, UltraFeedback (UF)), 2,000 preference samples were selected.
+Performance is highlighted as follows: bold numbers denote the best results, blue numbers indicate the significantly degraded results, and underlined numbers represent runner-up performances to the best number.
+P, Z, and N denote the most positive, near-zero, and most negative selection principles.
+Strategy Rand External Margin Implicit Margin IFD Margin BeeS Fullset
+Region P Z N P Z N P Z N P
+TL;DR 46.50 66.25 42.00 22.00 30.75 43.00 19.75 1.75 41.25 55.25 83.25 36.75
+HH 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 90.25 92.00
+UF 82.86 91.18 73.29 25.84 89.81 77.14 37.02 72.05 83.60 54.53 91.68 80.99
+We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare BeeS with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of BeeS in enhancing comprehensive dialogue ability in Section 4.3.
+Lastly, we perform diverse ablation studies for the BeeS in Section 4.4.
 4.1 Experimental Setup
-Table 1: GPT4 win rates vs.
-chosen answers on three tasks.
-DPO is implemented on subsets selected using different filtering strategies.
-The supervised fintuned model of Llama-3.2-3B-base is used for initialization.
-Each strategy selected 2,000 samples for DPO training across TL;DR, HH, and Ultrafeedback.
-Bold and underlined numbers indicate the best and runner-up performances, respectively.
-Filter Init Rand External Margin Implicit Margin IFD Margin DM Full
-Strategy TOP MID BOT TOP MID BOT TOP MID BOT ADD MUL
-TL;DR 34.00 46.50 66.25 42.00 22.00 30.75 43.00 19.75 1.75 41.25 55.25 63.75 83.25 36.75
-HH 70.25 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 85.25 87.25 92.00
-Ultrafeedback 61.49 82.86 91.18 73.29 25.84 89.81 77.14 37.02 72.05 83.60 54.53 91.55 90.68 80.99
 Preference Datasets.
-We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (Bai et al., 2022), and (3) UltraFeedback (Cui et al., 2024), which comprises quality-scored model responses across diverse prompts from multiple sources.
-To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (Meng et al., 2024).
+We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (volske2017tl, ; stiennon2020learning, ) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (bai2022training, ), and (3) UltraFeedback (cui2023ultrafeedback, ), which comprises quality-scored model responses across diverse prompts from multiple sources.
+To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (meng2024simpo, ).
 In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model.
-The scores are given by the PairRM (Jiang et al., 2023b) reward model.
+The scores are given by the PairRM (jiang2023llm, ) reward model.
 Statistics about these datasets are in Appendix A.
 Models.
-We perform sample selection and DPO training experiments across three model architectures: Llama-3.2-3B (MetaAI, 2024), Llama-3-8B (Dubey et al., 2024), and Mistral-7B (Jiang et al., 2023a), under Base and Instruct setup.
+Our experiments are conducted across four model series: Llama-3.2 (llama32, ), Llama-3 (dubey2024llama, ), Mistral-7B-v2 (jiang2023mistral, ), and Qwen-2.5 (yang2024qwen2, ) under Base and Instruct setups.
 For the Base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental Instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets.
-For the instruct setup, we directly use Llama-3-8B-Instruct or Mistral-7B-Instruct-v0.2 as the start of DPO training.
-Concerning external reward model for margin calculation, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) which is the best reward model at this scale according to the RewardBench leadboard.
-Evaluation.
-Following Rafailov et al.
-(2024), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately.
-For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (Li et al., 2023b) as our evaluation benchmark, which consists of 805 diverse questions.
-[^1] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (Ye et al., 2023), whereas AlpacaEval 2.0 employs GPT4-1106-preview.
-As the ground truth oracle is unavailable, we assess model performance through win rates against reference responses (written by humans or LLMs like GPT4).
-Specifically, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue.
+For the Instruct setup, we directly use them as the start of DPO training.
+Regarding the external reward model, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (liu2024skywork, ) that is the best reward model at this scale according to the RewardBench leadboard.
+As for the implicit reward model, we employ the Llama-3.2-3B Base and its DPO-tuned model (on 2,000 randomly selected samples from the complete dataset) for $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$.
+Implementation and Evaluation.
+For DPO training, we follow (rafailov2024direct, ) and use a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$.
+We run each training for two epochs, with a learning rate of $5\times 10^{-7}$, and a $0.1$ warmup ratio.
+Following (rafailov2024direct, ), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately.
+For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (li2023alpacaeval, ) as our evaluation benchmark, which consists of 805 diverse questions.
+[^2] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (ye2023comprehensive, ), whereas AlpacaEval 2.0 employs GPT4-1106-preview As the ground truth oracle is unavailable, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue.
 We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments.
@@
 Baselines.
-For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (Xia et al., 2024)), (2) Conditional Perplexity (IFD) (Li et al., 2023a), which utilizes token probability to measure informativeness.
+For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (xia2024less, )), (2) IFD (li2023quantity, ) (i.e., exponential form of the Point-wise Mutual Information), which measures semantic overlap.
 We use the difference in IFD scores between chosen and rejected responses for preference data selection.
 (3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards.
-For (2) and (3), We segment the data into TOP (highest positive difference), MID (near-zero difference), and BOT (lowest negative difference) subsets.
-Specifically, previous work (Wu et al., 2024) argues that low-gap preference pairs are more informative (i.e., chosen and rejected answers are similar), here we use the IFD-MID strategy to quantify this scheme and call it Low-Gap.
-For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by Xiong et al.
-(2024); Dong et al.
-(2024) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
-Implementation Details.
-The SFT training of the base model is carried out for two epochs with a learning rate of $2\times 10^{-5}$.
-Sample packing (Tunstall et al., 2024) is employed to accelerate the training, and we use a block size of 4096.
-As for DPO training, we follow Rafailov et al.
-(2024) and used a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$.
-The DPO training process runs for two epochs using the AdamW optimizer.
-The learning rate follows a cosine schedule with a 0.1 warm-up ratio, and the global batch size is set to 32.
-For implicit reward calculation, we employ a Llama-3.2-3B SFT model and its DPO-tuned model, which is fine-tuned on 2,000 randomly selected samples from the complete dataset.
-More implementation details are presented in Appendix B.
+For (2) and (3), we segment the data into P (most positive pairs), Z (close to zero pairs), and N (most negative pairs) subsets according to margin values.
+Specifically, previous work (wu2024beta, ) posits that "hard" preference pairs (where chosen and rejected samples are highly similar) are more beneficial for training, and we use the IFD-Z to quantify this scheme and call it Low-Gap.
+For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by (xiong2024iterative, ; dong2024rlhf, ) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
+4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
+First, we compare BeeS and baseline strategies on three widely-used preference datasets: TL;DR, HH, and UltraFeedback.
+Using a Llama-3.2-3B model as our Base architecture, we evaluate different selection methods, each sampling 2,000 training examples for DPO training.
+We use AlpacaEval as the test sets of UltraFeedback as it better reveals the degree of improvement.
+The results, measured by GPT4 win rates, are presented in Table 1.
+We summarize the findings below.
+  • Our method, BeeS, consistently achieves optimal or near-optimal win rates across all evaluated tasks, while all baseline methods show weak performance on at least one task.
+This outcome highlights BeeS’s superior robustness to noisy or detrimental samples across diverse task environments.
+Further, more data in DPO training does not always yield better results.
+Using just 2-5% of carefully selected data can surpass the performance achieved with the full dataset.
+Insights from Table 1 also reveal the existence of toxic samples and potential pitfalls of certain selection strategies.
+For instance, results highlighted by blue numbers show that models trained on data selected using external, implicit, or IFD margins can sometimes perform significantly worse than models trained on randomly chosen subsets.
+Such outcomes highlight the critical need for rigorous data quality assessment and effective filtering mechanisms in DPO training pipelines.
+  • Among all methods, only implicit margin-N consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit reward margin into the BeeS strategy.
+Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences.
+As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
+4.3 AlpacaEval 2.0 Win Rate Comparison
 Table 2: Performance comparison on AlpacaEval 2.0 using DPO-trained models with different 6,000-sample subsets (10% of full set).
@@
 Bold number denotes the best-performing selected subset.
+Blue numbers denote results that show little advantage over random.
 Dataset UltraFeedback Llama-UltraFeedback
@@
 Init 9.61 6.69 22.92 22.57 9.61 6.69 22.92 22.57
-Full 17.32 15.30 28.64 26.54 19.92 16.45 32.31 32.44
 Random 12.33 10.96 22.74 24.59 11.58 9.51 31.51 31.92
@@
 M-Im 19.33 17.80 29.71 29.44 18.88 16.25 33.71 32.92
-DM-ADD 18.64 18.46 29.75 29.62 20.98 18.29 35.97 35.79
-DM-MUL 19.53 19.09 28.30 28.39 21.67 20.01 36.36 36.47
-4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
-Firstly, we compare DM and baseline strategies on three widely-used preference datasets: TL;DR, HH, and UltraFeedback.
-Using a Llama-3.2-3B model as our base architecture, we evaluate different selection methods, each sampling 2,000 training examples for DPO training.
-We use AlpacaEval as the test sets of UltraFeedback as it better reveals the degree of improvement.
-The results, measured by GPT4 win rates, are presented in Table 1.
-We summarize the findings below.
-  • DM consistently achieves (near) optimal win rates across all three tasks, while all baselines show weak results on at least one task.
-This demonstrates that our method is more robust to useless or toxic samples in diverse domains.
-For example, DM-MUL achieves 17% higher win rates to second most effective baselines on TL;DR task, this is because the DM-MUL assigns low selection priority so long as a low margin is observed from one source of reward.
-  • More data in DPO training does not always yield better results - using just 2-5% of carefully selected data can surpass the performance achieved with the full dataset.
-This finding suggests the presence of ineffective samples in these tasks, as evidenced by the performance patterns of X-MID and X-BOT strategies (where X represents implicit margin).
-The results in Table 1 reveal that external, implicit, and IFD margins sometimes can identify toxic samples, as models trained on such data perform significantly worse than the original SFT model.
-This phenomenon underscores the critical importance of data filtering in DPO training.
-highlighting that training data should undergo a thorough quality assessment before being used in the training process.
-  • Among all methods, only implicit margin-BOT consistently identifies toxic samples, emphasizing the value of incorporating DPO implicit rewards into the DM strategy design.
-Despite its strong performance in RewardBench, the Skywork reward model’s margin signals prove less effective than random selection on HH, highlighting the Out-of-Distribution challenge external reward models face when evaluating unfamiliar behavioral patterns/preferences.
-As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
-4.2.1 Analysis
-We analyze DM from two perspectives: the joint distribution of different sources of reward margins and the training dynamics for subsets with different margin properties.
-We utilize the UltraFeedback datasets for this analysis.
-The left and middle panels of Figure LABEL:fig:visual illustrate the sample distribution across external-implicit and implicit-implicit margin dimensions on UltraFeedback.
-Several key observations emerge: (1) The correlation between implicit and external reward margins is notably weak.
-In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa.
-This pattern is consistent across other datasets (see Appendix C.2).
-This divergence highlights the distinct preference patterns captured by these two reward types.
-Supporting this observation, Table 1 demonstrates that neither margin type consistently outperforms the other in subset selection: external margins prove more effective for TL;DR, implicit margins show superior performance on HH, while both perform comparably on UltraFeedback.
-These findings underscore the importance of combining both reward types in an ensemble approach.
-(2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B), and the robust correlation is observed on other datasets as well (see Appendix C.2).
-This validates our design of using a separate small model for implicit reward margin calculations.
-The right panel of Figure LABEL:fig:visual shows DPO training loss curves across different subset selection strategies.
-TOP/BOT selections converge quickly to low loss values, while MID selections maintain steady, higher loss levels.
-Surprisingly, BOT-selected samples learn as quickly as TOP-selected ones, suggesting that ‘bad’ preferences are also easy to grasp for LLM.
-The pattern is consistent across different datasets and models (see Appendix D.1 for more results concerning loss and margin curves on other datasets).
-While this observation might explain proposals that use absolute margin values for selection (Muldrew et al., 2024), Table 1 reveals that TOP and BOT samples ultimately produce opposing effects despite similar training dynamics.
-However, we observe a difference between BOT area samples selected by external and implicit margins, where Im-BOT exhibits training dynamics that are more similar to Im-TOP area samples in terms of loss reduction speed.
-4.3 AlpacaEval 2.0 Win Rate Comparison
-In this section, we conduct a detailed examination of the chat task to understand how data filtering enhances both DPO training efficiency and models’ versatile conversational abilities compared to GPT-4-turbo, representing a key application area for preference learning.
-We use both Llama-3-8B (base) and (instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, a benchmark widely adopted in preference learning studies.
-The results of DM and various baselines are shown in Table 2.
-Table 3: Iterative DPO: AlpacaEval 2.0 benchmark results across three iterations of DPO training with UltraFeedback prompts.
+BeeS 19.53 19.09 30.03 30.46 21.67 20.01 36.36 36.47
+Full 17.32 15.30 28.64 26.54 19.92 16.45 32.31 32.44
+In this section, we aim to understand how data filtering influences DPO training efficiency and models’ versatile conversational abilities, representing a key application area for preference learning.
+We use both Llama-3-8B (Base) and (Instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, and results are shown in Table 2, LABEL:fig:loss-margin-8b and 6.
+BeeS consistently outperforms fullest DPO training and other selection strategies.
+As shown in Table 2, BeeS-selected subsets achieve around 4% higher win rates compared to full dataset training across all four settings.
+This distinct advantage highlights BeeS’s superior data and training efficiency, and further confirms the significant value of effective data filtering for DPO training.
+In contrast, all baseline strategies demonstrate inferior performance or limited improvement on some evaluated settings (see blue results).
+We attribute this instability to samples with ambiguous preferences, and whose margins differ a lot for different reward models.
+Different training dynamics of ‘P’/‘N’/‘Z’ subset region.
+The left panel of Figure LABEL:fig:loss-margin-8b shows DPO training loss curves for subsets selected by various strategies.
+Notably, training on subsets filtered according to the ‘P’ and ‘N’ criteria results in a rapid decrease in loss.
+In contrast, the loss curves corresponding to the ‘Z’ criterion tend to stabilize at consistently higher plateaus.
+Notably, ’N’-selected samples, which are often assumed as "difficult-to-learn" (gao2025principled, ), can actually be learned as rapidly as ’P’-selected samples, suggesting that ‘bad’ preferences are also easy to grasp for LLM.
+The pattern is consistent across different datasets and models (see Appendix C.1 for more results.
+While this observation might explain proposals that use absolute margin values for selection (muldrew2024active, ), Table 1 reveals that ‘P’ and ‘N’ samples produce opposing effects despite similar training dynamics.
+The middle and right panels of Figure LABEL:fig:loss-margin-8b illustrate that data subsets selected by BeeS exhibit both the most rapid decrease in training loss and the fastest increase in the DPO training margin, i.e., current train-batch average implicit margin.
+These concurrent observations of accelerated optimization help to explain the superior performance achieved by BeeS.
 Method Llama-3-Base (8B) Llama-3-Instruct (8B)
@@
 DPO (r3) 29.03 30.86 2736 44.51 53.12 2860
-DPO-DM (r1) 16.35 13.09 1624 42.20 45.74 2091
-DPO-DM (r2) 23.79 24.17 1901 46.40 50.60 2316
-DPO-DM (r3) 32.31 33.91 2565 48.49 54.99 2774
-DM-ADD and DM-MUL consistently outperform full dataset DPO training.
-Table 2 shows models trained on DM-selected subsets (using only 10% of data) achieve 2-4% higher raw and LC win rates compared to full dataset training across different models (base and instruct) and datasets (offline and on-policy).
-These consistent improvements demonstrate DM’s robustness and effectiveness, validating the value of data filtering for DPO training.
-In contrast, all baseline methods demonstrate inconsistent performance across different datasets and models.
-We attribute this instability to ambiguous samples whose margins vary significantly across different reward models, suggesting that proper fusion of reward signals is necessary to mitigate such noise.
-Appendix D.2 presents additional training loss and margin curves, illustrating distinct training dynamics across different subsets, and the trend is similar to that on the 3B model as in Appendix D.1.
-On-policy data primarily benefits its corresponding on-policy model.
-Llama-3-8B-Instruct shows approximately 8% higher win rates when trained on the on-policy Llama-UltraFeedback data compared to the off-policy UltraFeedback data, while Llama-3-8B-base demonstrates notably smaller improvements between these datasets.
-This phenomenon persists in both full and subset DPO training.
-4.4 Online Iterative DPO
-We explore the benefits of data selection in iterative DPO training using prompts from UltraFeedback as in Xiong et al.
-(2024).
-Our methodology employs DM-MUL to filter 5,000 samples from a pool of 20,000 generated samples per iteration.
-We evaluate both the base and instruct versions of Llama-3-8B, conducting single-epoch training rounds with hyperparameters $\beta=0.1$ and learning rate $5\times 10^{-7}$.
-Table 3 presents our experimental findings, from which we draw two key observations:
-While Best-of-N and Worst-of-N preference pairs from a strong reward model are informative, data filtering is still important to DPO training.
-DPO-DM performs better for both base and instruct settings using only 25% of the data.
-This can be attributed to the presence of numerous ambiguous, low-margin samples in the constructed online preference dataset.
-Furthermore, standard iterative DPO shows increasing output length across training rounds, but data filtering mitigates this issue.
-By round three, DPO-DM demonstrates both superior performance and better output length control.
-Table 4: Ablation study on model choice (Mistral-7B-Instruct-v-0.2). Each strategy selects a 6,000-sample subset.
-Method LC (%) WR (%) Length
-Initial 17.11 14.72 1676
-Full 18.00 18.77 4189
-Random 19.42 16.64 1621
-Low-Gap 13.31 14.35 2525
-DM-ADD 23.67 18.21 1527
-DM-MUL 26.04 20.53 1568
+DPO-BeeS (r1) 16.35 13.09 1624 42.20 45.74 2091
+DPO-BeeS (r2) 23.79 24.17 1901 46.40 50.60 2316
+DPO-BeeS (r3) 32.31 33.91 2565 48.49 54.99 2774
+Figure 6: AlpacaEval 2.0 results for on-policy datasets: (Left) Iterative DPO results across three DPO training iterations using UltraFeedback prompts.
+(Right) DPO on Llama-UltraFeedback subsets of varying sizes, selected by BeeS.
+Results of DPO-variants trained on fullset are also compared.
+Extension to Iterative DPO.
+We explore the data efficiency of iterative DPO using prompts from UltraFeedback as in (xiong2024iterative, ).
+In comparison, 20k prompts are used for on-policy preference pair generation per iteration, and our online version uses BeeS to reserve only 5k samples per iteration.
+The results are in the left panel of Figure 6.
+There is high redundancy in the on-policy data construction manner.
+Although iterative DPO shows much higher data efficiency than one-pass DPO training (i.e., better results than those in Table 2), data selection is still important for quality control.
+This can be attributed to the presence of numerous ambiguous, low-margin samples (usually paired with low-quality prompts).
+A smaller $\beta$ value in DPO loss correlates with higher data efficiency.
+While $\beta$ is commonly recognized as a factor controlling the strength of the Kullback-Leibler (KL) divergence, it also significantly influences data efficiency.
+Specifically, the DPO loss, defined as $\log\sigma(\beta\times m_{im})$, indicates that a reduced $\beta$ allows for effective gradient updates for more preference pairs with wider margins.
+To investigate this, we conducted DPO training on the Llama-3-Instruct 8B model using its on-policy dataset, Llama-UltraFeedback, with $\beta$ set to $0.01$.
+We then evaluated performance using varying numbers of samples selected by our method, BeeS.
+The results, presented in the right panel of Figure 6, demonstrate that relaxing the margin constraint in the DPO loss substantially improves data efficiency (Refer to Appendix C.4 for results on the Base setup).
+Notably, DPO training with a 3k-sample subset selected by BeeS achieved performance comparable to training with the full dataset (which is 20 times larger).
+Furthermore, we compared BeeS data selection with several established DPO variants that modify the original loss function, including SimPO (meng2024simpo, ), $\beta$-DPO (wu2024beta, ), and $\alpha$-DPO (wu2024alpha, ).
+Our findings indicate that BeeS is unique in its ability to effectively enhance both the win rate and the length-controlled (LC) win rate.
+In contrast, these variants primarily improved the LC win rate, and to a lesser extent than BeeS.
+These results underscore the significant potential of data selection and data efficiency considerations to enhance the original DPO training algorithm.
 4.4 Ablation Study
-A critical aspect of dataset filtering methods is their generalization capability - specifically, how well the method (selected subset) transfers to new models, similar optimization algorithms, etc.
-From this perspective, we examine DM’s robustness and effectiveness across several dimensions.
-Data filtering remains effective for new architecture and datasets.
-While our initial experiments focused on the Llama architecture and tokenizers, we investigated DM’s effectiveness when applied to fundamentally different model architectures.
-We evaluate the Mistral-7B-Instruct model using its corresponding Mistral-UltraFeedback dataset, maintaining consistent experimental parameters with those described in Section 4.3.
-The AlpacaEval 2.0 evaluation results are presented in Table 4.
-Our experiments show that DM-MUL significantly outperforms baseline methods while maintaining output lengths (i.e.
-around 1600) comparable to the initial instruct model.
-We observe that models trained on the full dataset tend to generate notably longer responses (i.e., 4189), primarily due to their learned behavior of incorporating numerous hyperlinks (e.g., https://www.ibdb.com/) in an attempt to enhance response credibility.
-Similarly, models trained on the Low-Gap selected subset exhibit increased response lengths (i.e., 2525) and follow a comparable pattern of including hyperlinks.
-In contrast, models trained on DM-selected subsets do not display this behavior.
-We consider the inclusion of hyperlinks without proper RAG (Retrieval-Augmented Generation) (Lewis et al., 2020) or website verification mechanisms to be problematic, as such references can be unreliable.
-DM selected subsets are effective across other preference learning algorithms.
-We examine the effectiveness of our selected subsets across different DPO variants and compare the raw/LC win rates of models trained on full or selected subsets by IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024) and SLiC (Zhao et al., 2023).
-The results are presented in Figure LABEL:fig:abl_po.
-Our results demonstrate that DM-M consistently outperforms full-set training across all three algorithms, achieving particularly notable improvements in raw/LC win rates - over 12% - for the IPO algorithm.
+A critical aspect of dataset filtering methods is their generalization capability—specifically, the performance when transferred to new models or applied with similar optimization algorithms.
+Data filtering remains effective for new LLM architecture.
+we evaluated BeeS on several contemporary model architectures: Mistral-7B-Instruct, Qwen-2.5-7B-Instruct, and Qwen-2.5-14B-Instruct.
+Consistent with previous experiments, BeeS was used to select a 10% data subset, and its performance was compared against DPO training on the full dataset.
+The AlpacaEval 2.0 evaluation results are presented in the left panel of Figure 7.
+We observe that BeeS consistently and significantly outperforms full-dataset DPO training.
+Notably, even though larger models like Qwen-14B inherently demonstrate higher data efficiency, our data selection strategy, BeeS, still improved the win rate by approximately 3% while utilizing only 10% of the data.
+BeeS selected subsets are effective for diverse preference learning algorithms.
+We examine whether the subset selected by BeeS remains data efficient for DPO-variants like IPO (azar2024general, ), KTO (ethayarajh2024kto, ), and SLiC (zhao2023slic, ).
+We utilize the BeeS selected 6k-sample subset from Llama-UltraFeedback and the results are presented in the right panel of Figure 7.
+We observe that the high-margin subset consistently benefit these preference learning algorithms by outperforming full-set training.
+Notably, it achieves large improvements in raw/LC win rates — over 12% for the IPO algorithm.
 This advantage is maintained even across these preference learning algorithms with varying data efficiency (as measured by the performance gap between randomly selected 6,000 samples and the full dataset).
 These findings highlight the significant value of sample filtering for other preference learning.
-Additional results related to instruct model training can be found in Appendix D.4.
-[…]
-5 Related Work
-Preference Learning Algorithms.Reinforcement Learning from Human Feedback has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (Ouyang et al., 2022).
-While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed.
-These include but not limited other RL-based training algorithms (Li et al., 2023c; Zhong et al., 2024), rejection sampling (Dong et al., 2023; Gulcehre et al., 2023), conditional supervised fine-tuning (Lu et al., 2022; Yang et al., 2024; Zhang et al., 2024a), and Direct Preference Optimization (Rafailov et al., 2024).
-Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance.
-Following the introduction of DPO, numerous works (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024; Tang et al., 2024; Han et al., 2024; Xu et al., 2024; Hong et al., 2024; Park et al., 2024) have attempted to improve its performance by modifying the DPO objective.
-Data Selection in LLM Fine-tuning.
-Data selection is crucial in LLM post-training (Wang et al., 2024), primarily due to two key observations: LLM training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (Swayamdipta et al., 2020; Deng et al., 2023).
-Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (Cao et al., 2024; Li et al., 2024; Xia et al., 2024), often adapting active learning query strategies (Ren et al., 2021) to assess sample uncertainty and diversity.
-However, data efficiency in preference learning remains relatively unexplored.
-While some studies studied annotation cost reduction in preference dataset creation (Muldrew et al., 2024; Yasunaga et al., 2024), our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance.
-Furthermore, our method extends to iterative DPO (Xiong et al., 2024) and its variants (Zhang et al., 2024b), where training data is generated online.
+Additional results related to Instruct model training can be found in Appendix C.3.
+Mistral-7B Qwen-2.5-7B Qwen-2.5-14B
+LC (%) WR (%) LC (%) WR (%) LC (%) WR (%)
+Initial 17.11 14.72 31.27 31.60 37.03 32.64
+Full 18.00 18.77 39.67 38.24 49.99 46.81
+BeeS 26.04 20.53 46.20 43.78 50.20 49.74
+Figure 7: Ablation Study: (Left) different model choices (Mistral-7B-Instruct-v-0.2, Qwen-2.5-Instruct-7B and Qwen-2.5-Instruct-14B).
+BeeS selects a 6k-sample subset for training.
+(Right) variants of DPO: win rate comparison on IPO, KTO, and SLiC algorithms.
+UltraFeedback is used for the preference learning on Llama-3-8B (Base) model.
+Rand and BeeS select a 6k-sample subset.
 5 Conclusion
 Our research bridges the gap between algorithmic advances and data-focused approaches in Large Language Model (LLM) alignment by systematically examining how preference data quality affects Direct Preference Optimization (DPO).
-We address the issue of parameter shrinkage caused by noisy data and introduce a dual-margin-based strategy for selecting high-quality training examples.
+We address the issue of parameter shrinkage caused by noisy data and introduce a BeeS strategy for selecting high-quality training examples.
 This approach not only improves model performance but also significantly increases computational efficiency.
@@
 Win rate judgments serve as a key evaluation metric across our experiments.
-For the TL;DR and HH, we slightly modify the evaluation prompts suggested by Rafailov et al.
-(2024) for the win rates judge.
+For the TL;DR and HH, we slightly modify the evaluation prompts suggested by [38] for the win rates judge.
 We use the original prompt designed for AlpacaEval and AlpacaEval 2.0 benchmark in all testing.
@@
 FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why.
-SECOND, on a new line, state only ”A” or ”B” to indicate your choice.
-Your response should use the format: Comparison: ¡one-sentence comparison and explanation¿ Preferred: ¡”A” or ”B”¿
+SECOND, on a new line, state only "A" or "B" to indicate your choice.
+Your response should use the format: Comparison: <one-sentence comparison and explanation> Preferred: <"A" or "B">
 {spverbatim}
@@
 FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless.
-SECOND, on a new line, state only ”A” or ”B” to indicate which response is more helpful or harmless.
-Your response should use the format: Comparison: ¡one-sentence comparison and explanation¿ More helpful or harmless: ¡”A” or ”B”¿
-Appendix B More Implementation Details
-We present the implementation details of our baseline methods: TOP, MID, and BOT using margins calculated from three metrics, Conditional Perplexity (CPPL), External (Ex), and Implicit (Im) rewards.
-Subsequently, we describe the implementation of DM-MUL in our experimental setup.
+SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful or harmless.
+Your response should use the format: Comparison: <one-sentence comparison and explanation> More helpful or harmless: <"A" or "B">
+A.1 More Implementation Details
+SFT.
+The SFT training of the Base model is carried out for two epochs with a learning rate of $2\times 10^{-5}$.
+Sample packing [45] is employed to accelerate the training, and we use a block size of 4096.
+We present the implementation details of our baseline methods: P, Z, and N using margins calculated from three metrics, IFD/Conditional Perplexity (CPPL), External (Ex), and Implicit (Im) rewards.
+Subsequently, we describe the implementation of BeeS in our experimental setup.
 The baseline strategies are implemented as follows: first, we eliminate outlier samples with extreme margin values (both positively high and negatively low) for CPPL, Ex, and Im metrics.
-For the TOP and BOT strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively.
-The MID strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
-For DM-MUL, we set $M_{1}=-2$ as the lower bound for both external and implicit reward margins.
-The upper bound $M_{2}$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[M_{2},\text{max margin}]$ is less than 30, or (2) The number of samples in $[M_{2},\text{max margin}]$ is less than max $\text{max margin}-M_{2}$.
-Appendix C Visualization of Margin Distributions
-C.1 Singular Margin Distribution
+For the P and N strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively.
+The Z strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
+For BeeS, we set $L=-2$ as the lower bound for both external and implicit reward margins.
+The upper bound $U$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[U,\text{max margin}]$ is less than 30, or (2) The number of samples in $[U,\text{max margin}]$ is less than max $\text{max margin}-U$.
+Appendix B Visualization of Margin Distributions
+B.1 Singular Margin Distribution
 The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures LABEL:fig:cppl-margin,LABEL:fig:ex-margin,LABEL:fig:im-margin, reveal a notable concentration of sample margins around zero.
 This clustering around the zero indicates ambiguous preference labels.
-It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 4.2.1.
-C.2 Joint Margin Distribution
+It leads to the challenge in preference learning, as evidenced by the substantially slower decrease in training loss (and slower increase in training margin) compared to samples with larger margins, as shown in Section 3.2.
+B.2 Joint Margin Distribution
 To complement the left and middle subfigures in Figure LABEL:fig:visual, we present additional results showing the joint margin distributions of samples on the other datasets in Figure LABEL:fig:visual-extra.
 Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation.
-These further enhance the rationality of our design detail of DM: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
-Appendix D More Experimental Results
-D.1 Train Loss and Margin Curves - 3B
+These further enhance the rationality of our design detail of BeeS: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
+Appendix C More Experimental Results
+C.1 Train Loss and Margin Curves - 3B
 To complement the right subfigure in Figure LABEL:fig:visual, we present additional results showing the progression of training loss and margins throughout the DPO training process.
 The results are shown in Figure LABEL:fig:loss-margin.
-All strategies demonstrated consistent patterns in loss reduction: both TOP margin-oriented and BOT strategies achieved rapid decreases in training loss, while the MID strategy exhibited slower convergence and remained at significantly higher final loss values.
-Regarding training margins, TOP strategies achieved higher levels compared to BOT and MID approaches.
-Notably, our proposed DM-ADD and DM-MUL strategies demonstrated even larger margins than the Implicit Margin-TOP strategy.
-D.2 Train Loss and Margin Curves - 8B
-Figure LABEL:fig:loss-margin-8b presents the progression of training loss and margins during DPO training of the 8B model.
-We show the fine-tuned version of the Llama-3-8B (base) model trained on UltraFeedback as a representative case since all models exhibited similar trends, as in Figure LABEL:fig:loss-margin.
-D.3 Resources and computation cost
+All strategies demonstrated consistent patterns in loss reduction: both P margin-oriented and N strategies achieved rapid decreases in training loss, while the Z strategy exhibited slower convergence and remained at significantly higher final loss values.
+Regarding training margins, P strategies achieved higher levels compared to N and Z approaches.
+Notably, our proposed BeeS strategy demonstrated even larger margins than the Implicit Margin-P strategy.
+C.2 Resources and computation cost
 For all experiments, we utilized 8 A100 GPUs.
@@
 In each round of iterative DPO implementation, we performed generation and annotation processes on 4 A100 GPUs, with each GPU processing 5,000 prompts with 5 distinct generations per prompt.
-The overall generation that utilizes vLLM (Kwon et al., 2023) for acceleration takes about 1.5 hours, and the corresponding reward annotation takes about 2 hours.
-D.4 More Results for Ablation Study on the DPO Variants
-As a complementary study to the results shown in Figure LABEL:fig:abl_po, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters.
-The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by DM-MUL achieved significantly higher win rates across most evaluation scenarios.
-D.5 Ablation Study on Hyper-parameters
-We investigate the impact of hyperparameters $\beta$ and training epochs on DPO performance.
-Specifically, we train the Llama-UltraFeedback model using DPO with $\beta=0.01$ for both one and two epochs.
-The model’s performance is evaluated using AlpacaEval 2.0 win rates, with results shown in Table 6.
-Our experimental results demonstrate that DM-MUL selected subsets significantly enhance the performance of vanilla DPO on these different hyperparameter configurations.
-Notably, DPO-DM-MUL achieves superior results compared to state-of-the-art preference learning algorithms such as SimPO (Meng et al., 2024), despite the latter utilizing the complete dataset for DPO training.
-Table 6: Ablation study on $\beta$ ($\beta=0.01$): performance comparison of methods trained on Llama-Ultrafeedback dataset, benchmarked against state-of-the-art algorithms trained on the complete dataset.
-Method LC (%) WR (%) Epochs
-DPO 38.65 39.57 2
-DPO 40.30 37.90 1
-DPO-DM-MUL 43.66 42.52 2
-DPO-DM-MUL 45.17 41.86 1
-KTO 33.10 31.80 1
-IPO 35.60 35.60 1
-SimPO 43.72 39.88 1
-However, we observe that selecting a small $\beta$ value is not consistently a good choice for DPO training.
+The overall generation that utilizes vLLM [23] for acceleration takes about 1.5 hours, and the corresponding reward annotation takes about 2 hours.
+C.3 More Results for Ablation Study on the DPO Variants
+As a complementary study to the results shown in Figure 7, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters.
+The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by BeeS achieved significantly higher win rates across most evaluation scenarios.
+C.4 Hyperparameters Risks
+Although smaller $\beta$ can lead to higher data efficiency, we observe that it can bring potential issues for DPO training.
 Specifically, a small $\beta$ corresponds to a relaxed Kullback-Leibler (KL) divergence constraint in the policy optimization process.
 This relaxation can permit excessive deviation from the initial policy, potentially compromising the model’s learned behaviors and stability during training.
-For instance, when we conduct DPO training with Llama-3-8B-Instruct/Mistral-7B-Instruct-v0.2 on the Ex/Im-TOP selected 6,000 subsets from UltraFeedback, with $\beta=0.01$ and two epochs update, we find that although the model could respond normally to most questions, it sometimes outputs repeated or chaotic tokens, as shown in Figure 9.
+For instance, when we conduct DPO training with Llama-3-8B-Instruct/Mistral-7B-Instruct-v0.2 on the Ex/Im-P selected 6,000 subsets from UltraFeedback, with $\beta=0.01$ and two epochs update, we find that although the model could respond normally to most questions, it sometimes outputs repeated or chaotic tokens, as shown in Figure 10.
 And their win rates on AlpacaEval 2.0 dropped by more than 10 points as a consequence.
@@
 These operations can lead to a relatively smaller drop in chosen/rejected log probability.
-However, we still do not recommend using small $\beta$, as more potential sacrifices in other tasks such as reasoning may also appear in this case.
 Figure 10: The model breaking pattern when conducting DPO training with small $\beta$ ($\beta=0.01$) for Llama-3-8B-Instruct and Mistral-7B-Instruct-V0.2.
 We select two examples of abnormal responses given by each model.
+Appendix D Limitations.
+The empirical evaluations in this study primarily focused on models up to the 14B parameter scale, where BeeS demonstrated notable efficacy.
+Extending these investigations to significantly larger foundation models, such as those in the 70B parameter range or beyond, was constrained by the computational resources available for the current work.
+Future research could build upon our findings by exploring the scalability and performance of BeeS in these larger-scale settings.

### v4

@@
 1 Introduction
-Reinforcement Learning from Human Feedback (RLHF; christiano2017deep, ; ziegler2019fine, ) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values.
+Reinforcement Learning from Human Feedback (RLHF; Christiano et al., 2017; Ziegler et al., 2019) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values.
 Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization.
 However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
-As an alternative, (rafailov2024direct, ) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data.
+As an alternative, (Rafailov et al., 2024) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data.
 DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF.
-Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ) and iterative learning schemes (xiong2024iterative, ).
+Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024) and iterative learning schemes (Xiong et al., 2024).
 While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
@@
 Preference learning algorithms.
-Reinforcement Learning from Human Feedback also known as dueling RL (pacchiano2021dueling, ) or preference-based RL (chen2022human, ), has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (ouyang2022training, ).
+Reinforcement Learning from Human Feedback also known as dueling RL (Pacchiano et al., 2021) or preference-based RL (Chen et al., 2022), has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (Ouyang et al., 2022).
 While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed.
-These include but not limited other RL-based training algorithms (li2023remax, ; zhong2024dpo, ), rejection sampling (dong2023raft, ; gulcehre2023reinforced, ), conditional supervised fine-tuning (lu2022quark, ; yang2024rewards, ; zhang2024reward, ), and Direct Preference Optimization (rafailov2024direct, ).
+These include but not limited other RL-based training algorithms (Li et al., 2023c; Zhong et al., 2024), rejection sampling (Dong et al., 2023; Gulcehre et al., 2023), conditional supervised fine-tuning (Lu et al., 2022; Yang et al., 2024b; Zhang et al., 2024a), and Direct Preference Optimization (Rafailov et al., 2024).
 Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance.
-Following the introduction of DPO, numerous works (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ; meng2024simpo, ; tang2024generalized, ; han2024f, ; xu2024contrastive, ; hong2024orpo, ; wu2024self, ) have attempted to improve its performance by modifying the DPO objective.
+Following the introduction of DPO, numerous works (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024; Tang et al., 2024; Han et al., 2024; Xu et al., 2024; Hong et al., 2024; Wu et al., 2024c) have attempted to improve its performance by modifying the DPO objective.
 Data selection in LLM Fine-tuning.
-Data selection is crucial in LLM post-training (wang2024survey, ) for two key observations: post-training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (swayamdipta2020dataset, ; deng2023counterfactual, ).
-Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (cao2023instruction, ; li2024superfiltering, ; xia2024less, ), often adapting active learning query strategies (ren2021survey, ) to assess sample uncertainty and diversity.
+Data selection is crucial in LLM post-training (Wang et al., 2024) for two key observations: post-training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (Swayamdipta et al., 2020; Deng et al., 2023b).
+Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (Cao et al., 2024; Li et al., 2024; Xia et al., 2024), often adapting active learning query strategies (Ren et al., 2021) to assess sample uncertainty and diversity.
 However, data efficiency in preference learning remains relatively unexplored.
-Prior studies have studied reducing annotation costs in preference dataset creation (muldrew2024active, ; yasunaga2024alma, ) and on scenarios involving numerous ranking annotations (thekumparampil2024comparing, ; mukherjee2024optimal, ).
-Other research aims to improve a model’s ability to distinguish between two responses by adding a margin to the loss term (meng2024simpo, ; park2024disentangling, ; azar2024general, ).
-Additionally, concurrent work highlights the importance of margins for preference data filtering, though there is debate on whether hard samples help or hinder preference learning (wu2024beta, ; khaki2024rs, ; yu2025rip, ; gao2025principled, ).
+Prior studies have studied reducing annotation costs in preference dataset creation (Muldrew et al., 2024; Yasunaga et al., 2024) and on scenarios involving numerous ranking annotations (Thekumparampil et al., 2024; Mukherjee et al., 2024).
+Other research aims to improve a model’s ability to distinguish between two responses by adding a margin to the loss term (Meng et al., 2024; Park et al., 2024; Azar et al., 2024).
+Additionally, concurrent work highlights the importance of margins for preference data filtering, though there is debate on whether hard samples help or hinder preference learning (Wu et al., 2024b; Khaki et al., 2024; Yu et al., 2025; Gao et al., 2025).
 Our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance.
-Furthermore, our method extends to iterative DPO (xiong2024iterative, ) and its variants (zhang2024self, ), wherein training data is dynamically generated by the model during its iterative training process.
+Furthermore, our method extends to iterative DPO (Xiong et al., 2024) and its variants (Zhang et al., 2024b), wherein training data is dynamically generated by the model during its iterative training process.
 2 Background
@@
 In the reward learning stage, a reward model is trained to approximate human preferences based on preference data.
-By adopting the Bradley-Terry model (bradley1952rank, ) to capture human preference, reward training involves minimizing the loss:
+By adopting the Bradley-Terry model (Bradley and Terry, 1952) to capture human preference, reward training involves minimizing the loss:
 $\mathcal{L}_{\mathrm{RM}}(r)=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\big[\log\sigma\big(r(x,y_{w})-r(x,y_{l})\big)\big],$
@@
 Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model.
-This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; schulman2017proximal, ) to optimize the policy by maximizing the expected reward.
+This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; Schulman et al., 2017) to optimize the policy by maximizing the expected reward.
 $\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[r(x,y)-\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}\right],$
 where $\beta>0$ is the regularization parameter. However, this RL approach can be computationally expensive, sensitive to reward misspecification and require careful hyperparameter tuning.
-Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; rafailov2024direct, ) has been proposed.
+Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; Rafailov et al., 2024) has been proposed.
 DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training.
@@
 3.1 Parameter Shrinkage and Inflation Analysis
-We follow the model from (zhu2023principled, ) to illustrate why data selection can improve model performance.
+Figure 1: The workflow of the BeeS method.
+We follow the model from (Zhu et al., 2023) to illustrate why data selection can improve model performance.
 We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$.
-For reward learning, our reward model can be an explicit $r(x,y)$ (ouyang2022training, ), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (rafailov2024direct, ).
+For reward learning, our reward model can be an explicit $r(x,y)$ (Ouyang et al., 2022), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (Rafailov et al., 2024).
 Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model.
@@
 $\mathbb{E}_{x,\zeta}\Big[\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big]\leq\mathbb{E}_{x}\Big[\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big].$
-Since the optimal $\omega$ is $\omega^{*}$ without $\zeta$, plugging $\omega^{*}$ in Equation 2 will cause the left-hand side to be greater than the right-hand side.
+Since the optimal $\omega$ is $\omega^{*}$ without $\zeta$, plugging $\omega^{*}$ in Equation˜2 will cause the left-hand side to be greater than the right-hand side.
 Therefore, the optimal $\omega$ with the existence of $\zeta$ intends to shrink to the original point compared to $\omega^{*}$ so that the first-order condition is still satisfied.
 We provide the underlying intuition with an extreme example.
-If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Equation 1.
+If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Equation˜1.
 In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero.
@@
 Finally, we remark that $\zeta$ can come from multiple resources, including human classification errors, different embeddings or reward models from other LLMs and so on.
+Table 1: Symbols used in the formulation.
+r x $y_{w}/y_{l}$ $\phi/\Delta\phi$ $\zeta$ w
+Reward Input prompt Preferred/dispreferred Response (relative) feature function Exogenous error Learnable parameters
 Parameter Inflation. We next explain why selecting data points based on the margin can lead to parameter inflation, thereby offsetting the parameter shrinkage caused by errors.
@@
 Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side.
-Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Equation 1.
-Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Equation 1 continuously.
+Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Equation˜1.
+Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Equation˜1 continuously.
 Hence, some techniques like regularization or early stopping when training are indispensable.
@@
 3.2 Multi-source Margin Aggregation
+(a) Visualization of joint margin distribution on UltraFeedback.
+(Left) Joint distribution of external and implicit reward margin values.
+(Middle) Joint distribution of implicit reward margins computed using models of 1B and 3B scales.
+(Right) Joint distribution of two different external reward margin values on online-generated data.
 Building on our previous analysis, we aim to develop a data selection strategy based on the margin-maximization principle, with the calculation of reward margin being the critical component.
 We examine two distinct types of reward margin calculations: external reward margin and implicit reward margin.
-The external reward margin is provided by an external reward function, while the implicit reward margin is derived from the implicit reward $\log\frac{\pi_{\theta}}{\pi_{\mathrm{ref}}}$, where $\pi_{\theta}$ represents the policy trained by DPO (rafailov2024direct, ) (see Section 4.1 for details on margin calculation).
-We visualize the joint distribution of different reward margin sources using the UltraFeedback dataset (cui2023ultrafeedback, ) in Figure LABEL:fig:visual.
-  • The left and middle panels of Figure LABEL:fig:visual reveal several key phenomena: (1) The correlation between implicit and external reward margins is notably weak.
+The external reward margin is provided by an external reward function, while the implicit reward margin is derived from the implicit reward $\log\frac{\pi_{\theta}}{\pi_{\mathrm{ref}}}$, where $\pi_{\theta}$ represents the policy trained by DPO (Rafailov et al., 2024) (see Section 4.1 for details on margin calculation).
+We visualize the joint distribution of different reward margin sources using the UltraFeedback dataset (Cui et al., 2024) in Figure 4(a).
+  • The left and middle panels of Figure 4(a) reveal several key phenomena: (1) The correlation between implicit and external reward margins is notably weak.
 In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa.
@@
 Notably, these two patterns remain consistent across other datasets as well (see Appendix B.2).
-  • Online RLHF (xiong2024iterative, ; dong2025self, ) employs the target model to generate multiple responses for given prompts iteratively and uses an external reward model to identify the response pair with the largest margin for DPO training.
-The right panel of Figure LABEL:fig:visual illustrates that a max-margin pair construction method, even when derived from one strong reward model, can still yield ambiguous preferences when evaluated by another reward model (which shows similar performance on RewardBench Leadboard).
+  • Online RLHF (Xiong et al., 2024; Dong et al., 2025) employs the target model to generate multiple responses for given prompts iteratively and uses an external reward model to identify the response pair with the largest margin for DPO training.
+The right panel of Figure 4(a) illustrates that a max-margin pair construction method, even when derived from one strong reward model, can still yield ambiguous preferences when evaluated by another reward model (which shows similar performance on RewardBench Leadboard).
 This ambiguity in preference signal indicates that the online data generation process can still cause high redundancy, which may offer little to no benefit, or could even be detrimental, to online-DPO training.
@@
 Our objective is to obtain the in-distribution implicit reward model with low computational and GPU-memory cost.
-Given the strong correlation of that margin across different models (See Figure LABEL:fig:visual), this weak-to-strong guidance is feasible.
+Given the strong correlation of that margin across different models (See Figure 4(a)), this weak-to-strong guidance is feasible.
 To achieve this, we randomly select a small seed dataset $\mathcal{D}_{0}$ from $\mathcal{D}$ and employ a (or several) small model to perform preference learning on this seed set.
-The high sample and training efficiency of the DPO loss (kim2025spread, ) ensures the feasibility of this approach.
+The high sample and training efficiency of the DPO loss (Kim et al., 2025) ensures the feasibility of this approach.
 Step 2: margin calculation.
@@
 This projection serves to quantify the confidence that a specific preference direction $y_{w}>y_{l}$ is correct, formally expressed as $\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})$.
-Assuming that these sources are conditionally independent, then following the Bayesian formula transformation in previous work (liu2022mpc, ; deng2024a3s, ), the preference probability can be expressed as:
+Assuming that these sources are conditionally independent, then following the Bayesian formula transformation in previous work (Liu et al., 2022; Deng et al., 2024), the preference probability can be expressed as:
 $\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})=\frac{\prod_{i=1}^{K}\mathbb{P}(y_{w}>y_{l}|m^{i})}{\prod_{i=1}^{K}\mathbb{P}(y_{w}>y_{l}|m^{i})+\prod_{i=1}^{K}(1-\mathbb{P}(y_{w}>y_{l}|m^{i}))}.$ (3)
@@
 4 Experiments
+We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare BeeS with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of BeeS in enhancing comprehensive dialogue ability in Section 4.3.
+Lastly, we perform diverse ablation studies for the BeeS in Section 4.4.
+4.1 Experimental Setup
+Preference Datasets.
+We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (Bai et al., 2022), and (3) UltraFeedback (Cui et al., 2024), which comprises quality-scored model responses across diverse prompts from multiple sources.
+To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (Meng et al., 2024).
+In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model.
+The scores are given by the PairRM (Jiang et al., 2023b) reward model.
+Statistics about these datasets are in Appendix A.
+Models.
+Our experiments are conducted across four model series: Llama-3.2 (MetaAI, 2024), Llama-3 (Dubey et al., 2024), Mistral-7B-v2 (Jiang et al., 2023a), and Qwen-2.5 (Yang et al., 2024a) under Base and Instruct setups.
+For the Base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental Instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets.
+For the Instruct setup, we directly use them as the start of DPO training.
+Regarding the external reward model, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) that is the best reward model at this scale according to the RewardBench leadboard.
+As for the implicit reward model, we employ the Llama-3.2-3B Base and its DPO-tuned model (on 2,000 randomly selected samples from the complete dataset) for $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$.
+Implementation and Evaluation.
+For DPO training, we follow (Rafailov et al., 2024) and use a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$.
+We run each training for two epochs, with a learning rate of $5\times 10^{-7}$, and a $0.1$ warmup ratio.
+Following (Rafailov et al., 2024), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately.
+For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (Li et al., 2023b) as our evaluation benchmark, which consists of 805 diverse questions.
+[^2] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (Ye et al., 2023), whereas AlpacaEval 2.0 employs GPT4-1106-preview As the ground truth oracle is unavailable, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue.
+We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments.
+More details are presented in Appendix A.
+Baselines.
+For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (Xia et al., 2024)), (2) IFD (Li et al., 2023a) (i.e., exponential form of the Point-wise Mutual Information), which measures semantic overlap.
+We use the difference in IFD scores between chosen and rejected responses for preference data selection.
+(3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards.
+For (2) and (3), we segment the data into P (most positive pairs), Z (close to zero pairs), and N (most negative pairs) subsets according to margin values.
+Specifically, previous work (Wu et al., 2024b) posits that "hard" preference pairs (where chosen and rejected samples are highly similar) are more beneficial for training, and we use the IFD-Z to quantify this scheme and call it Low-Gap.
+For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by (Xiong et al., 2024; Dong et al., 2024) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
+We provide source code of our paper in https://github.com/xiangtanshi/DPO-Data-Selection.
+4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
 Table 2: GPT-4 judged win rates for Llama-3.2-3B models fine-tuned with DPO on subsets selected by various data selection strategies.
-For every strategy and benchmark (TL;DR, HH, UltraFeedback (UF)), 2,000 preference samples were selected.
+For every strategy and benchmark (TL;DR, $H$, UltraFeedback (UF)), 2,000 preference samples were selected.
 Performance is highlighted as follows: bold numbers denote the best results, blue numbers indicate the significantly degraded results, and underlined numbers represent runner-up performances to the best number.
@@
 TL;DR 46.50 66.25 42.00 22.00 30.75 43.00 19.75 1.75 41.25 55.25 83.25 36.75
-HH 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 90.25 92.00
+$H$ 84.25 82.25 76.50 69.75 92.25 81.25 32.00 11.25 90.00 64.50 90.25 92.00
 UF 82.86 91.18 73.29 25.84 89.81 77.14 37.02 72.05 83.60 54.53 91.68 80.99
-We organize the experiments as follows: we explain the experimental setup in Section 4.1; we compare BeeS with various sample selection baselines on diverse preference tasks and present the detailed results in Section 4.2; then we focus on the important chat task, and explore the effectiveness of BeeS in enhancing comprehensive dialogue ability in Section 4.3.
-Lastly, we perform diverse ablation studies for the BeeS in Section 4.4.
-4.1 Experimental Setup
-Preference Datasets.
-We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (volske2017tl, ; stiennon2020learning, ) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (bai2022training, ), and (3) UltraFeedback (cui2023ultrafeedback, ), which comprises quality-scored model responses across diverse prompts from multiple sources.
-To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (meng2024simpo, ).
-In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model.
-The scores are given by the PairRM (jiang2023llm, ) reward model.
-Statistics about these datasets are in Appendix A.
-Models.
-Our experiments are conducted across four model series: Llama-3.2 (llama32, ), Llama-3 (dubey2024llama, ), Mistral-7B-v2 (jiang2023mistral, ), and Qwen-2.5 (yang2024qwen2, ) under Base and Instruct setups.
-For the Base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental Instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets.
-For the Instruct setup, we directly use them as the start of DPO training.
-Regarding the external reward model, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (liu2024skywork, ) that is the best reward model at this scale according to the RewardBench leadboard.
-As for the implicit reward model, we employ the Llama-3.2-3B Base and its DPO-tuned model (on 2,000 randomly selected samples from the complete dataset) for $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$.
-Implementation and Evaluation.
-For DPO training, we follow (rafailov2024direct, ) and use a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$.
-We run each training for two epochs, with a learning rate of $5\times 10^{-7}$, and a $0.1$ warmup ratio.
-Following (rafailov2024direct, ), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately.
-For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (li2023alpacaeval, ) as our evaluation benchmark, which consists of 805 diverse questions.
-[^2] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (ye2023comprehensive, ), whereas AlpacaEval 2.0 employs GPT4-1106-preview As the ground truth oracle is unavailable, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue.
-We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments.
-More details are presented in Appendix A.
-Baselines.
-For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (xia2024less, )), (2) IFD (li2023quantity, ) (i.e., exponential form of the Point-wise Mutual Information), which measures semantic overlap.
-We use the difference in IFD scores between chosen and rejected responses for preference data selection.
-(3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards.
-For (2) and (3), we segment the data into P (most positive pairs), Z (close to zero pairs), and N (most negative pairs) subsets according to margin values.
-Specifically, previous work (wu2024beta, ) posits that "hard" preference pairs (where chosen and rejected samples are highly similar) are more beneficial for training, and we use the IFD-Z to quantify this scheme and call it Low-Gap.
-For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by (xiong2024iterative, ; dong2024rlhf, ) and run for three rounds, each using 20k prompts sampled from UltraFeedback.
-4.2 Win Rate Comparison with Baselines on Classic Preference Datasets
 First, we compare BeeS and baseline strategies on three widely-used preference datasets: TL;DR, HH, and UltraFeedback.
@@
 As for the IFD margin metric, it exhibits notable inconsistency across different datasets, rendering it unreliable for evaluating new datasets where prior preference patterns are unknown.
+In general, preference learning departs from traditional representation learning, which predominantly leverages contrastive samples to improve discriminative capacity (Deng et al., 2023b, a).
+Preference learning focuses on capturing semantic relationships, and models benefit when the underlying preferences in the data are explicit and well-defined.
 4.3 AlpacaEval 2.0 Win Rate Comparison
@@
 Full 17.32 15.30 28.64 26.54 19.92 16.45 32.31 32.44
+(a) DPO training loss and margin of Llama-3.2-3B Base (Left) and Llama-3-8B Base (Middle and Right) on UltraFeedback datasets.
 In this section, we aim to understand how data filtering influences DPO training efficiency and models’ versatile conversational abilities, representing a key application area for preference learning.
-We use both Llama-3-8B (Base) and (Instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, and results are shown in Table 2, LABEL:fig:loss-margin-8b and 6.
+We use both Llama-3-8B (Base) and (Instruct) models, measuring performance through raw and length-controlled win rates on AlpacaEval 2.0, and results are shown in Table 3, Figures˜7(a) and 7.
 BeeS consistently outperforms fullest DPO training and other selection strategies.
@@
 Different training dynamics of ‘P’/‘N’/‘Z’ subset region.
-The left panel of Figure LABEL:fig:loss-margin-8b shows DPO training loss curves for subsets selected by various strategies.
+The left panel of Figure 7(a) shows DPO training loss curves for subsets selected by various strategies.
 Notably, training on subsets filtered according to the ‘P’ and ‘N’ criteria results in a rapid decrease in loss.
 In contrast, the loss curves corresponding to the ‘Z’ criterion tend to stabilize at consistently higher plateaus.
-Notably, ’N’-selected samples, which are often assumed as "difficult-to-learn" (gao2025principled, ), can actually be learned as rapidly as ’P’-selected samples, suggesting that ‘bad’ preferences are also easy to grasp for LLM.
+Notably, ’N’-selected samples, which are often assumed as "difficult-to-learn" (Gao et al., 2025), can actually be learned as rapidly as ’P’-selected samples, suggesting that ‘bad’ preferences are also easy to grasp for LLM.
 The pattern is consistent across different datasets and models (see Appendix C.1 for more results.
-While this observation might explain proposals that use absolute margin values for selection (muldrew2024active, ), Table 1 reveals that ‘P’ and ‘N’ samples produce opposing effects despite similar training dynamics.
-The middle and right panels of Figure LABEL:fig:loss-margin-8b illustrate that data subsets selected by BeeS exhibit both the most rapid decrease in training loss and the fastest increase in the DPO training margin, i.e., current train-batch average implicit margin.
+While this observation might explain proposals that use absolute margin values for selection (Muldrew et al., 2024), Table 2 reveals that ‘P’ and ‘N’ samples produce opposing effects despite similar training dynamics.
+The middle and right panels of Figure 7(a) illustrate that data subsets selected by BeeS exhibit both the most rapid decrease in training loss and the fastest increase in the DPO training margin, i.e., current train-batch average implicit margin.
 These concurrent observations of accelerated optimization help to explain the superior performance achieved by BeeS.
@@
 Extension to Iterative DPO.
-We explore the data efficiency of iterative DPO using prompts from UltraFeedback as in (xiong2024iterative, ).
+We explore the data efficiency of iterative DPO using prompts from UltraFeedback as in (Xiong et al., 2024).
 In comparison, 20k prompts are used for on-policy preference pair generation per iteration, and our online version uses BeeS to reserve only 5k samples per iteration.
@@
 Notably, DPO training with a 3k-sample subset selected by BeeS achieved performance comparable to training with the full dataset (which is 20 times larger).
-Furthermore, we compared BeeS data selection with several established DPO variants that modify the original loss function, including SimPO (meng2024simpo, ), $\beta$-DPO (wu2024beta, ), and $\alpha$-DPO (wu2024alpha, ).
+Furthermore, we compared BeeS data selection with several established DPO variants that modify the original loss function, including SimPO (Meng et al., 2024), $\beta$-DPO (Wu et al., 2024b), and $\alpha$-DPO (Wu et al., 2024a).
 Our findings indicate that BeeS is unique in its ability to effectively enhance both the win rate and the length-controlled (LC) win rate.
@@
 BeeS selected subsets are effective for diverse preference learning algorithms.
-We examine whether the subset selected by BeeS remains data efficient for DPO-variants like IPO (azar2024general, ), KTO (ethayarajh2024kto, ), and SLiC (zhao2023slic, ).
+We examine whether the subset selected by BeeS remains data efficient for DPO-variants like IPO (Azar et al., 2024), KTO (Ethayarajh et al., 2024), and SLiC (Zhao et al., 2023).
 We utilize the BeeS selected 6k-sample subset from Llama-UltraFeedback and the results are presented in the right panel of Figure 8.
@@
 These results emphasize the importance of careful data curation in developing better alignment techniques and provide practical guidelines for future research and implementation.
+Acknowledgements
+This work is supported by the National Natural Science Foundation of China (U24B20180) and the Alibaba Group through Alibaba Research Intern Program.
+We appreciate the reviewers for their insightful feedback and advice, these constructive criticism and recommendations have been invaluable in helping us improve the quality of this work.
 Appendix A Datasets and Evaluation Details
@@
     TL;DR     92858     400     Summarization
-            HH             86372             400             Helpful & Harmless
+    $H$     86372     400     Helpful & Harmless
     UltraFeedback     61135     -     Hybrid
@@
 We use the original prompt designed for AlpacaEval and AlpacaEval 2.0 benchmark in all testing.
-{spverbatim}
 **TL;DR win rate prompt**: Which of the following summaries does a better job of summarizing the most important points in the given forum post, without including unimportant or irrelevant details?
 A good summary is both precise and concise.
-Post: Post
-Summary A: Summary_A
-Summary B: Summary_B
+Post:
+{Post}
+Summary A:
+{Summary_A}
+Summary B:
+{Summary_B}
 FIRST provide a one-sentence comparison of the two summaries, explaining which you prefer and why.
 SECOND, on a new line, state only "A" or "B" to indicate your choice.
-Your response should use the format: Comparison: <one-sentence comparison and explanation> Preferred: <"A" or "B">
-{spverbatim}
+Your response should use the format:
+Comparison: <one-sentence comparison and explanation>
+Preferred: <"A" or "B">
 **HH win rate prompt**: For the following query to a chatbot, which response is more helpful or harmless?
@@
 Query: {query}
-Response A: res_a
-Response B: res_b
+Response A:
+{res_a}
+Response B:
+{res_b}
 FIRST provide a one-sentence comparison of the two responses and explain which you feel is more helpful or harmless.
 SECOND, on a new line, state only "A" or "B" to indicate which response is more helpful or harmless.
-Your response should use the format: Comparison: <one-sentence comparison and explanation> More helpful or harmless: <"A" or "B">
+Your response should use the format:
+Comparison: <one-sentence comparison and explanation>
+More helpful or harmless: <"A" or "B">
 A.1 More Implementation Details
@@
 For the P and N strategies, we select samples based on their ranking positions at the upper and lower ends of the distribution, respectively.
-The Z strategy involves random sampling from the subset of samples whose margin values fall within the interval $[-\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
-For BeeS, we set $L=-2$ as the lower bound for both external and implicit reward margins.
-The upper bound $U$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $[U,\text{max margin}]$ is less than 30, or (2) The number of samples in $[U,\text{max margin}]$ is less than max $\text{max margin}-U$.
+The Z strategy involves random sampling from the subset of samples whose margin values fall within the interval $\tau,\tau]$, where $\tau$ is set to 0.1 for CPPL and 1.0 for Ex/Im metrics.
+For BeeS, we set L=2L=-2 as the lower bound for both external and implicit reward margins.
+The upper bound $U$ is determined dynamically based on two conditions: (1) The number of samples with margin values in the interval $U,\text{max margin}]$ is less than 30, or (2) The number of samples in $U,\text{max margin}]$ is less than $\text{max margin}-U$.
+Figure 9: The model breaking pattern when conducting DPO training with small $\beta$ ($\beta=0.01$) for Llama-3-8B-Instruct and Mistral-7B-Instruct-V0.2.
+We select two examples of abnormal responses given by each model.
 Appendix B Visualization of Margin Distributions
+(a) Distribution of implicit reward margins on TL;DR, HH, and UltraFeedback datasets.
+The reward is calculated using the Llama-3.2-3B SFT model, and its weakly aligned DPO model that is fine-tuned on 2,000 randomly selected samples from the full set.
+(b) Distribution of external rewards on TL;DR, HH, and UltraFeedback datasets. The reward is calculated using Skywork-Reward-Llama-3.1-8B-v0.2.
+(c) Distribution of conditional perplexity (also named instruction following difficulty) margins on TL;DR, HH, and UltraFeedback datasets. The perplexity is calculated with the Llama-3.2-3B SFT model.
 B.1 Singular Margin Distribution
-The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures LABEL:fig:cppl-margin,LABEL:fig:ex-margin,LABEL:fig:im-margin, reveal a notable concentration of sample margins around zero.
+The margin distributions calculated using CPPL margin, External and Implicit DPO reward margins, as illustrated in Figures 12(c),12(b),12(a), reveal a notable concentration of sample margins around zero.
 This clustering around the zero indicates ambiguous preference labels.
@@
 B.2 Joint Margin Distribution
-To complement the left and middle subfigures in Figure LABEL:fig:visual, we present additional results showing the joint margin distributions of samples on the other datasets in Figure LABEL:fig:visual-extra.
+To complement the left and middle subfigures in Figure 4(a), we present additional results showing the joint margin distributions of samples on the other datasets in Figure 12(d).
 Our analysis reveals that external and implicit margins exhibit minimal correlation across all four datasets, while implicit margins calculated by different models maintain a high correlation.
 These further enhance the rationality of our design detail of BeeS: fusion of both margins and disentangling implicit margin from the target model (if the target model is a bit large and we want to accelerate the enumeration process of the full-set.)
+(d) Subfigure (a)-(d): scatter plot showing the joint distribution of samples across external and implicit reward margin values on four datasets.
+Subfigure (e)-(f): joint distribution of implicit reward margins computed using models of 1B and 3B scales on two datasets.
 Appendix C More Experimental Results
 C.1 Train Loss and Margin Curves - 3B
-To complement the right subfigure in Figure LABEL:fig:visual, we present additional results showing the progression of training loss and margins throughout the DPO training process.
-The results are shown in Figure LABEL:fig:loss-margin.
+To complement the right subfigure in Figure 4(a), we present additional results showing the progression of training loss and margins throughout the DPO training process.
+The results are shown in Figure 12(e).
 All strategies demonstrated consistent patterns in loss reduction: both P margin-oriented and N strategies achieved rapid decreases in training loss, while the Z strategy exhibited slower convergence and remained at significantly higher final loss values.
@@
 Notably, our proposed BeeS strategy demonstrated even larger margins than the Implicit Margin-P strategy.
+(e) DPO train loss and margin on TL;DR, HH, and UltraFeedback datasets. The training was implemented with Llama-3.2-3B SFT version on different subsets selected by five strategies.
 C.2 Resources and computation cost
@@
 As a complementary study to the results shown in Figure 8, we conducted experiments using the Llama-3-8B-Instruct model while maintaining all other experimental parameters.
-The results, presented in Figure LABEL:fig:abl_po_ins, demonstrate that models trained on subsets selected by BeeS achieved significantly higher win rates across most evaluation scenarios.
+The results, presented in Figure 12(f), demonstrate that models trained on subsets selected by BeeS achieved significantly higher win rates across most evaluation scenarios.
+(f) Ablation study on variants of DPO: win rate comparison on IPO, KTO, and SLiC algorithms.
+The experiments utilize the UltraFeedback dataset for preference optimization, with the Llama-3-8B-Instruct model as the initial model.
+Random and BeeS select 6,000 samples (10% of the full set) for subset training.
 C.4 Hyperparameters Risks
+(g) DPO on Llama-UltraFeedback subsets of varying sizes, selected by BeeS. The training is conducted on Llama-3-8B-Base model.
 Although smaller $\beta$ can lead to higher data efficiency, we observe that it can bring potential issues for DPO training.
@@
 These operations can lead to a relatively smaller drop in chosen/rejected log probability.
-Figure 10: The model breaking pattern when conducting DPO training with small $\beta$ ($\beta=0.01$) for Llama-3-8B-Instruct and Mistral-7B-Instruct-V0.2.
-We select two examples of abnormal responses given by each model.
 Appendix D Limitations.
@@
 Future research could build upon our findings by exploring the scalability and performance of BeeS in these larger-scale settings.
+NeurIPS Paper Checklist
+The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
+Do not remove the checklist: The papers not including the checklist will be desk rejected.
+The checklist should follow the references and follow the (optional) supplemental material.
+The checklist does NOT count towards the page limit.
+Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
+  • You should answer [Yes] , [No] , or [N/A] .
+  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
+  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
+The checklist answers are an integral part of your paper submission.
+They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
+You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
+The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
+While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
+In general, answering "[No] " or "[N/A] " is not grounds for rejection.
+While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
+All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
+If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
+IMPORTANT, please:
+  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
+  • Keep the checklist subsection headings, questions/answers and guidelines below.
+  • Do not modify the questions and only use the provided macros for your answers.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Our abstract and introduction accurately reflect the paper’s contributions on data selection for direct preference learning.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: See Appendix D.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provide the derivation in Section 3.1.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: We provide all implementation details in the experimental part and appendix.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: All the datasets and models used in this work are publicly available.
+We utilize the open-source TRL repo for all our DPO experiments, which is easy to implement.
+All the details required to reproduce the main experimental results can be found in the experiments and appendix, and the code is available at https://github.com/xiangtanshi/DPO-Data-Selection.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We provide all training and test details in experimental section and appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: LLM experiments are typically costly and relatively robust, so we do not conduct the experiments repeatedly or report statistical significance metrics.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Refer to Appendix C.2
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Yes, the research in this paper fully conforms with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: [N/A]
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: [N/A]
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Yes, the paper properly credits the original creators of all assets used, explicitly mentions licenses and terms of use, and respects these conditions throughout the research.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: [N/A]
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: [N/A]
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: [N/A]
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: [N/A]
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.